### v1

Abstract
Accurate covariance forecasting is central to portfolio allocation, risk management, and asset pricing, yet many existing methods struggle at medium-term horizons, where shifting market regimes and slower dynamics predominate. We propose a deep learning framework that combines three-dimensional convolutional neural networks, bidirectional long short-term memory layers, and multi-head attention to capture complex spatio-temporal dependencies. Using daily data on 14 exchange-traded funds from 2017 through 2023, we find that our model reduces Euclidean and Frobenius distance metrics by up to 20% relative to classical benchmarks (e.g., shrinkage and GARCH approaches) and remains robust across distinct market regimes. Our portfolio experiments demonstrate significant economic value through lower volatility and moderate turnover. These findings highlight the potential of advanced deep learning architectures to improve medium-term covariance forecasts, offering practical benefits for institutional investors and risk managers.
Keywords: Covariance Matrix Forecasting, Deep Learning, Portfolio Optimisation, Medium-Term Horizons, Financial Markets.
JEL codes: C45, G11, G17.
1 Introduction
Covariance matrix estimation is central in modern finance. It influences portfolio construction, risk management, and asset pricing models. Markowitz, (1952) introduced mean-variance optimisation, highlighting how correlation among assets shapes the risk-return trade-off. Beyond portfolio choice, covariance estimation underpins factor-based models, hedging strategies, and even debates on market efficiency, where accurate correlation forecasts can drive risk control and potential outperformance.
Early studies found that even simple correlation structures could improve capital allocation. For instance, Elton and Gruber, (1973) suggested using constant-correlation estimates over historical windows for a decade. Their work showed that stable or slowly changing correlation patterns can help avoid distortions caused by transient market fluctuations.
Over time, multiple research strands have focused on short-horizon covariance forecasts. High-frequency datasets, such as intraday prices, enable detailed volatility analyses, often using generalised autoregressive conditional heteroskedasticity (GARCH) based models. Constant (Bollerslev,, 1990) and dynamic conditional correlation (Engle,, 2002) specifications remain common benchmarks, although Symitsi et al., (2018) argue that little fundamental progress has been made recently. Extensions include applying the heterogeneous autoregressive model (Corsi,, 2009) in a multivariate setting through Cholesky decomposition [^1] Cholesky decomposition is a method that breaks down a positive definite matrix into the product of a lower triangular matrix and its transpose. (Chiriac and Voev,, 2011). However, the Cholesky decomposition models outperform DRD-based [^2] Dynamic Relationship/Dependence (DRD) models refer to time-varying models that separate covariance into volatility and correlation components forecasts. (Bollerslev et al.,, 2018).
However, GARCH-type approaches struggle with large cross-sections, falling prey to the curse of dimensionality (Robert F. Engle and Wolf,, 2019). Tools like principal component analysis and random matrix theory (Ledoit and Wolf,, 2003; Laloux et al.,, 2000) have been effective for large datasets. Recent advances, such as dynamic conditional correlation with nonlinear shrinkage (Robert F. Engle and Wolf,, 2019), perform well overall, as per De Nard et al., (2024). Still, most of these methods focus on very short-term horizons, often just one step ahead or one day ahead, and rely on high-frequency data.
Long-horizon forecasts matter for institutional investors like pension funds and endowments, which rebalance portfolios over weeks or months. Moreover, high-frequency data are not always public or cost-effective, making the short-horizon focus less relevant for these institutions. Furthermore, slow-moving trends, such as equity–bond correlation shifts, can be overlooked if one relies solely on high-frequency econometric models. Sandoval and Franca, (2012) highlights how regime shifts or significant structural changes often go undetected in classical specifications, raising the risk of large drawdowns and poor capital allocation.
For longer-term horizons, naive, shrinkage, and factor models are more common (Ledoit and Wolf,, 2004; De Nard et al.,, 2022). Naive approaches assume that recent covariance estimates strongly predict future covariances as they follow a Markov process. At the same time, shrinkage techniques blend the sample covariance matrix with structured targets to reduce noise, particularly when the cross-section of assets is large (Ledoit and Wolf,, 2022). Factor models can provide more insight but may not generalise well across diverse portfolios, especially when markets are affected by many different factors. These methods are typically stable but can underperform when major shifts occur.
In recent years, there has been a surge in machine learning methods that can handle non-linearities and high dimensionality. Gu et al., (2020) and Kim et al., (2023) use techniques like random forests, support vector machines, and generative adversarial networks, while Zhang et al., (2024) applies graph neural networks to capture volatility spillovers for covariance matrix forecast. However, most of these solutions still emphasise short-term horizons.
Attention mechanisms (Vaswani et al.,, 2017) offer a promising direction by weighting key observations more heavily. Some studies apply these techniques in finance but still focus on short-horizon tasks (Nazareth and Ramana Reddy,, 2023; Olorunnimbe and Viktor,, 2023). In principle, attention-based models could capture the slow-moving, cyclical factors and abrupt shifts that define medium- and long-horizon covariance forecasts, realising this potential. However, it requires integrating sophisticated data-driven architectures with the economic logic of excess returns.
Recent efforts to extend these approaches to multi-asset portfolios have been limited. While many studies handle large cross-sections of stocks or focus on single-asset classes such as equities (Reis et al.,, 2024), fewer examine the combined dynamics across multiple asset classes in an integrated framework. However, institutional investors routinely diversify across equities, bonds, alternatives, and derivatives, underscoring the need for robust, multi-asset covariance forecasting models. In particular, capturing time-varying cross-asset correlations over medium-term horizons is a non-trivial challenge that remains under-explored in much of the existing literature.
Two key research gaps remain. First, most existing studies focus on short- or long-term horizons, leaving a relative void in medium-term covariance forecasting. This gap is particularly salient for multi-asset portfolios, where cross-asset correlations over multi-week or multi-month windows can significantly influence asset allocation and risk management (De Nard et al.,, 2021). Second, there is limited empirical evidence on whether using raw returns versus excess returns leads to systematically different covariance forecasts, especially over these medium-term horizons. Economic theory suggests that risk premiums and the underlying drivers of asset co-movements could be more accurately captured through excess returns. However, most machine learning and econometric approaches mostly default to raw returns.
Our study addresses these gaps by focusing on medium-term covariance forecasts in a multi-asset setting. Specifically, we propose a novel deep learning (DL) model that captures spatio-temporal correlations while remaining robust to structural changes. We benchmark against classical methods (naive, shrinkage, GARCH, and dimensional reduction) and demonstrate consistent gains in forecast accuracy under diverse market environments from 2017 to 2023. Our main contributions are as follows:
  • We present an integrated 3D-convolution neural network (CNN) and bidirectional long-short-term memory (LSTM) architecture with multi-head attention that improves medium-term forecasting accuracy.
  • We compare different return-processing specifications (including the treatment or omission of risk-free rates) and show that the proposed approach preserves its predictive edge.
  • We evaluate the economic value of our forecasts in a global minimum-variance (GMV) portfolio, demonstrating significant variance reduction and stable turnover.
The rest of the paper is structured as follows. Section 2 introduces the benchmark methods and our DL model. Section 3 presents the main empirical findings. Section 4 reports robustness checks. Section 5 explores implications for portfolio management. Finally, Section 6 concludes.
2 Methodology
Let $\mathbf{r}_{t}$ denote the $N$-dimensional row vector of excess daily returns for $N$ assets at time $t$. Assuming time-varying mean excess returns, $\mathbf{r}_{t}$ can be expressed as follows:
$\mathbf{r}_{t}=\bm{\mu}_{t}+\mathbf{e}_{t}$ (1)
where $\bm{\mu}_{t}$ is the $N$-dimensional vector of rolling means of the excess daily returns, and $\mathbf{e}_{t}$ is the i.i.d. error terms with a normal distribution within each rolling window:
$\mathbf{e}_{t}\sim N\left(\mathbf{0},\bm{\Sigma}_{t}\right)$ (2)
Although empirical evidence suggests financial returns often exhibit heavy tails and non-zero skewness (Cont,, 2001), we adopt the normality assumption because it is standard in many benchmark models (e.g., GARCH-based correlation models and Ledoit–Wolf shrinkage). However, we acknowledge that alternative distributions (e.g., Student-$t$ innovations, mixture models, or those explicitly modelling tail dependence) may better capture extreme events and skewness in practice [^3] Extending our DL framework to accommodate such heavy-tailed or asymmetric distributions is an area for future research. Nevertheless, our empirical results in subsequent sections suggest that even a normality-based design can significantly enhance covariance forecasts, given the robustness of the data-driven layers in handling complex, nonlinear dependencies. (Hansen,, 1994; Harvey,, 1997).
The rolling mean $\bm{\mu}_{t}$ and rolling standard deviation $\bm{\sigma}_{t}$ at time $t$ are calculated using a window of size $F$ as follows:
$\bm{\mu}_{t}=\frac{1}{F}\sum_{k=t-F+1}^{t}\mathbf{r}_{k}$ (3)
$\bm{\sigma}_{t}^{2}=\frac{1}{F-1}\sum_{k=t-F+1}^{t}\left(\mathbf{r}_{k}-\bm{\mu}_{t}\right)\odot\left(\mathbf{r}_{k}-\bm{\mu}_{t}\right)$ (4)
[…]
$Cov_{t}\left(\mathbf{r}_{i},\mathbf{r}_{j}\right)=\frac{1}{F-1}\sum_{k=t-F+1}^{t}\left(r_{i,k}-\mu_{i,\,t}\right)\left(r_{j,k}-\mu_{j,\,t}\right)$ (5)
Our goal is to estimate the conditional covariance matrix of the asset returns for the next $F$ days based on the information available at time $t$:
${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}=Cov\left(\mathbf{r}_{t+1},\mathbf{r}_{t+2},\ldots,\mathbf{r}_{t+F}\middle|\mathcal{F}_{t}\right)$ (6)
where $\mathcal{F}_{t}$ is the sigma-algebra representing all the information available at time $t$.
To estimate ${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}$, we compare naive and other benchmark models found in the literature alongside our proposed model. A description of each model is shown below.
2.1 Naïve ($NA$)
A naive covariance forecasting approach is based on the lagged realised covariance. This model assumes that covariance is a Markov process, so the covariance matrix of the previous period is highly informative about the future covariance matrix. Under this model:
${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}=\mathbf{\Sigma}_{t-F:t}$ (7)
This forecasting method is as simple as possible as it requires no optimisation.
2.2 Naïve full sample($NA^{F}$)
Another naïve technique is to use the full sample instead of the rolling covariance matrix as our estimator.
${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}=\mathbf{\Sigma}$ (8)
where $\mathbf{\Sigma}$ is the covariance matrix of our sample up until $t$.
2.3 Exponential Weighted Moving Average ($EWMA$)
The EWMA employs exponentially decaying weights for the covariance matrix. The covariance matrix in the EWMA model is recursively computed as follows:
${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}=(1-\eta)\mathbf{e}_{t}\mathbf{e}_{t}^{{}^{\prime}}+\eta\mathbf{\Sigma}_{t-F:t}$ (9)
where $\mathbf{e}_{t}$ is the N-dimensional vector of the error terms of all assets from Eq. (1), at time $t$, and $\eta$ acts as a decay factor, confined within the range [0,1], dictating the rate at which the weights on past observations decrease. This factor has been approximated to be around 0.94 (J.P.Morgan,, 1996).
2.4 Ledoit and Wolf ($LW$)
The LW shrinkage (Ledoit and Wolf,, 2004) is a robust statistical model that combines the sample covariance matrix with a structured target matrix, improving estimation accuracy by reducing the impact of estimation error due to high dimensionality and limited sample size. The LW shrinkage estimator is given by:
${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}=\rho\mathbf{T}+(1-\rho)\mathbf{\Sigma}_{t-F:t}$ (10)
where $\mathbf{T}=\frac{\text{Tr}(\mathbf{\Sigma}_{t-F:t})}{N}\times\mathbf{I}$ is the target matrix, with $\mathbf{I}$ being the identity matrix and $\text{Tr}(\mathbf{\Sigma}_{t-F:t})$ representing the sum of the diagonal elements of the sample covariance matrix. The shrinkage coefficient $\rho$ is optimally determined to minimise the mean squared error between the estimated and true covariance matrices:
$\rho=\frac{\sum_{i\neq j}\operatorname{Var}(\sigma_{i,j})}{\sum_{i\neq j}(\sigma_{i,j}-t_{i,j})}$ (11)
where $\sigma_{i,j}$ are the elements of the sample covariance matrix and $t_{i,j}$ are the elements of the target matrix $\mathbf{T}$.
2.5 Ledoit and Wolf full sample ($LW^{F}$)
Similarly to the $NA^{F}$ model, we employed a full sample of Ledoit and Wolf, (2004) as a shrinkage model.
${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}=\rho\mathbf{T}+(1-\rho)\mathbf{\Sigma}$ (12)
2.6 Constant Conditional Correlations ($CCC$)
The $CCC$ model, introduced by Bollerslev, (1990), assumes that asset correlations remain constant over time while variances evolve dynamically according to a GARCH process based on a one-step forecast. However, since investment funds and risk managers typically rebalance portfolios over extended horizons, such as monthly or quarterly, a one-step forecast is insufficient for practical decision-making (De Nard et al.,, 2022).
To extend the model to a multi-step forecasting framework, we follow the methodology of Baillie and Bollerslev, (1992), defining the $F$-step ahead conditional covariance forecast as:
$\widehat{\mathbf{\Sigma}}_{t+1:t+F}=\frac{1}{F}\sum_{f=1}^{F}\mathbf{D}_{t+f}\mathbf{R}\mathbf{D}_{t+f},$ (13)
where $\mathbf{D}_{t+f}=\text{diag}\left(\sqrt{h_{11,t+f}},\sqrt{h_{22,t+f}},\dots,\sqrt{h_{NN,t+f}}\right)$ is the diagonal matrix of forecasted conditional volatilities for horizon $f$ and $R$ the conditional correlation matrix. The multi-step averaging approach used in Eq. (13) was introduced by De Nard et al., (2021) to provide a more stable covariance forecast over the rebalancing period. The $h_{ii,t}$ series is modelled using univariate GARCH(1,1) processes:
$h_{ii,t+f}=\sum_{j=0}^{f-1}\omega_{i}(\alpha_{i}+\beta_{i})^{j}+(\alpha_{i}+\beta_{i})^{f}h_{ii,t},$ (14)
where $\omega_{i},\alpha_{i},\beta_{i}$ are the estimated GARCH parameters.
The conditional correlation matrix $\mathbf{R}$ is assumed to be constant over time calculated through the standardised residuals, which are defined as:
$z_{i,t}=\frac{e_{i,t}}{\sqrt{h_{ii,t}}},$ (15)
where $e_{i,t}$ are the residuals of the return series, which are used to estimate the conditional correlation matrix.
2.7 Dynamic Conditional Correlations ($DCC$)
The $DCC$ model of Engle, (2002) extends the $CCC$ model by allowing correlations to vary over time while retaining the GARCH framework for variances. This is achieved by replacing the constant correlation matrix in Eq. (13) with a dynamic correlation matrix $\mathbf{R}_{t}$.
To accommodate multi-step forecasting for practical portfolio management, we follow the methodology of Engle and Sheppard, (2001). The $F$-step ahead conditional covariance forecast is:
$\widehat{\mathbf{\Sigma}}_{t+1:t+F}=\frac{1}{F}\sum_{f=1}^{F}\mathbf{D}_{t+f}\mathbf{R}_{t+f}\mathbf{D}_{t+f},$ (16)
where $\mathbf{D}_{t+f}$ is defined as in Eq. (13), and $\mathbf{R}_{t+f}$ is the $f$-step ahead forecast of the dynamic conditional correlation matrix. The multi-step forecast for the correlation matrix is computed as follows:
$\mathbf{R}_{t+f}=\sum_{j=0}^{f-1}(1-\alpha-\beta)\overline{\mathbf{Q}}(\alpha+\beta)^{j}+(\alpha+\beta)^{f}\mathbf{R}_{t},$ (17)
where $\overline{\mathbf{Q}}$ is the unconditional correlation matrix, and $\alpha+\beta<1$ ensures stationarity. The parameters $\alpha$ and $\beta$ control the persistence of the dynamic correlations.
The conditional correlation matrix $\mathbf{R}_{t}$ evolves over time according to:
$\mathbf{R}_{t}=\mathbf{V}_{t}^{-1}\mathbf{Q}_{t}\mathbf{V}_{t}^{-1}$ (18)
$\mathbf{Q}_{t}=(1-\alpha-\beta)\overline{\mathbf{Q}}+\alpha\mathbf{z}_{t-1}\mathbf{z}_{t-1}^{\prime}+\beta\mathbf{Q}_{t-1},$ (19)
where $\mathbf{V}_{t}=\text{diag}\{\sqrt{q_{11,t}},\sqrt{q_{22,t}},\dots,\sqrt{q_{NN,t}}\}$ ensures that $\mathbf{R}_{t}$ remains a proper correlation matrix. The standardised residuals $\mathbf{z}_{t}$ are defined as in Eq. (15), and $\overline{\mathbf{Q}}$ is estimated from historical standardised residuals.
2.8 Nonlinear Shrinkage Dynamic Conditional Correlations ($DCC^{NL}$)
One recent extension is the $DCC^{NL}$ model from De Nard et al., (2022). The covariance forecast in the $DCC^{NL}$ model follows the same structure as in Eq. (16), but now with a shrinkage-improved correlation matrix ($\overline{\mathbf{Q}}$). They apply the nonlinear shrinkage method of Ledoit and Wolf, (2020), which optimally shrinks the eigenvalues of the sample covariance matrix towards a better-conditioned target matrix. This procedure ensures that the estimated correlation matrix remains positive, definite, and well-conditioned, even in high-dimensional settings.
Given a sample correlation matrix $\mathbf{\Upsilon}$, the optimal nonlinear shrinkage estimator adjusts its eigenvalues as follows. Let $\mathbf{\Upsilon}$ have the eigendecomposition:
$\mathbf{\Upsilon}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top},$ (20)
where $\mathbf{U}$ is the matrix of eigenvectors and $\mathbf{\Lambda}=\text{diag}(\lambda_{1},\lambda_{2},...,\lambda_{N})$ is the diagonal matrix of eigenvalues. The nonlinear shrinkage transformation replaces $\mathbf{\Lambda}$ with a shrinkage-transformed version $\mathbf{\Lambda}^{*}$:
$\mathbf{\Upsilon}^{*}=\mathbf{U}\mathbf{\Lambda}^{*}\mathbf{U}^{\top}$ (21)
where the optimally shrunk eigenvalues $\lambda_{i}^{*}$ are computed using the oracle nonlinear shrinkage function:
$\lambda_{i}^{*}=\lambda_{i}\frac{1}{\left(\pi c\lambda_{i}f(\lambda_{i})\right)^{2}+\left(1-c-\pi c\lambda_{i}H_{f}(\lambda_{i})\right)^{2}}$ (22)
where $f(\lambda_{i})$ is the sample spectral density estimator, $H_{f}(\lambda_{i})$ is the Hilbert transform of the spectral density, and $c=\frac{N}{T}$ is the limiting concentration ratio of the covariance matrix.
The sample spectral density function $f(x)$ is estimated using a kernel-based density estimator, ensuring uniform consistency. The Hilbert transform $H_{f}(x)$ is computed using the principal value integral, which corrects for eigenvalue bias induced by sampling variability.
The shrunk correlation matrix ($\mathbf{\Upsilon}^{*}$) is then used as the input for the $DCC$ process in Eq. (18).
2.9 Principal Component Analysis ($PCA$)
As noted by Ledoit and Wolf, (2003), empirical results show the usefulness of PCA in covariance matrix estimation. The PCA is computed as follows:
${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}=\mathbf{U\Lambda}\mathbf{U}^{\top}$ (23)
To ensure we capture most of the information, we retain only the first $k$ principal components that account for at least 95% of the total variance. This is mathematically expressed as:
$\sum_{i=1}^{k}\lambda_{i}\geq 0.95\sum_{i=1}^{N}\lambda_{i}$ (24)
The final forecasted covariance matrix is then approximated as:
${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}\approx\mathbf{U}_{k}\mathbf{\Lambda}_{k}\mathbf{U}_{k}^{\top}$ (25)
2.10 Random Matrix Theory ($RMT$)
$RMT$ (Laloux et al.,, 2000) is built upon the $PCA$ where the upper ($\varsigma_{+}$) and lower ($\varsigma_{-}$) bound of the eigenvalues are calculated using the Marchenko-Pastur distribution. These bounds are calculated as:
$\lambda_{+}=\left(1+\frac{1}{\sqrt{q}}\right)^{2},\quad\lambda_{-}=\left(1-\frac{1}{\sqrt{q}}\right)^{2}$ (26)
where $q=T/N$ is the ratio of the window size to the number of assets. Eigenvalues within these bounds are dominated by noise, while eigenvalues outside the bounds represent meaningful information.
$\lambda_{i}^{\text{filtered}}=\begin{cases}\lambda_{i}&\text{if }\lambda_{i}>\lambda_{+}\text{ or }\lambda_{i}<\lambda_{-}\\ \frac{\lambda_{+}+\lambda_{-}}{2}&\text{otherwise}\end{cases}$ (27)
This filtering process removes noise and retains only the significant components of the covariance matrix. The cleaned covariance matrix is reconstructed by multiplying the filtered eigenvalues with their corresponding eigenvectors:
${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}=\mathbf{U}\mathbf{\Lambda}^{\text{filtered}}\mathbf{U}^{\top}$ (28)
2.11 CNN with A-BiLSTM ($CAB$)
In our proposed model, each batch of input data is processed sequentially through all six stages, as described below.
2.11.1 Data Preprocessing
Let $\mathbf{B}=[\mathbf{D}_{1},\mathbf{D}_{2},\dots,\mathbf{D}_{T}]^{\top}$ be the time series data, where each $\mathbf{D}_{t}\in\mathbb{R}^{N\times N}$ represents the rolling covariance matrix, normalised, of $N$ assets at time $t$).
It is essential to standardise the input features to facilitate efficient training and convergence of the $CAB$ model. This transformation ensures that each feature has zero mean and unit variance, which is crucial for stabilising the training dynamics of neural networks.
We transform the data into overlapping sequences using a rolling window approach to capture temporal dependencies. Specifically, for a given lookback period $L$, the sequence creation process, at time $t$, can be expressed as:
$\mathbf{S}_{t}=[\mathbf{D}_{t-L},\mathbf{D}_{t-L+1},\dots,\mathbf{D}_{t}]^{\top}\in\mathbb{R}^{L\times N\times N}$ (29)
The input to the $CAB$ model consists of sequences of scaled covariance matrices. Formally, let $\mathbf{S}_{i}\in\mathbb{R}^{L\times N\times N}$ denote the $i$-th sequence.
2.11.2 3D Convolutional Layer
Next, each sequence of covariance matrices is processed through a three-dimensional layer (Lecun et al.,, 1998). We use a 3D CNN to capture local spatio-temporal features, given that each rolling covariance matrix is an ($N\times N$) grid, and the temporal dimension arises from stacking these grids over time. The 3D CNN filters can identify how subsets of assets interact and evolve over multiple past windows, offering a more flexible approach than linear assumptions.
By applying a convolution operation with a kernel size of ($ks\times ks\times ks$), the same padding, and a stride equal to one, the CNN layer captures patterns within the data across different time steps and different dimensions of the covariance matrices. The transformation performed by the 3D convolutional layer on the input sequence of covariance matrices can be expressed as:
$\mathbf{X}_{\text{conv}}(t,i,j)=\sum_{u=-\lfloor ks/2\rfloor}^{\lfloor ks/2\rfloor}\sum_{v=-\lfloor k_{s}/2\rfloor}^{\lfloor ks/2\rfloor}\sum_{w=-\lfloor ks/2\rfloor}^{\lfloor ks/2\rfloor}\mathbf{W}(u,v,w)\cdot\mathbf{X}_{\text{rs}}(t+u,i+v,j+w)+b$ (30)
where $ks$ is the kernel size of the 3D Convolution, determining the dimensions of the receptive field across the temporal and spatial axes. $\mathbf{W}(u,v,w)$ represents the kernel weights, a learnable parameter of size $C_{\text{out}}\times C_{\text{in}}\times k_{s}\times k_{s}\times k_{s}$, where $C_{\text{in}}$ is the number of input channels, and $C_{\text{out}}$ is the number of output channels. The input covariance matrices are reshaped to include a single channel, so $C_{\text{in}}=1$. Additionally, the model outputs a single channel, meaning $C_{\text{out}}=1$. Consequently, the weight tensor simplifies to $\mathbf{W}\in\mathbb{R}^{1\times 1\times ks\times ks\times ks}$. $\mathbf{X}_{\text{rs}}(t+u,i+v,j+w)$ is the input sequence of covariance matrices reshaped to dimensions $1\times L\times N\times N$. The term $b$ denotes the bias added to the convolution result, which, with $C_{\text{out}}=1$, is a single scalar, i.e., $b\in\mathbb{R}^{1}$. The indices $u$, $v$, and $w$ iterate over all whole integers within the range $[-(k_{s}/2),(k_{s}/2)]$, centered around $(t,i,j)$, to compute local features. Finally, zero-padding is applied to ensure that the output dimensions match the input dimensions, preserving the temporal and spatial resolution of the covariance matrices.
Subsequently, the convolutional output is flattened ($\mathbf{X}_{\text{flattened}}\in\mathbb{R}^{L\times N^{2}}$) to transition from spatial to feature dimensions suitable for the LSTM layer. This reshaping consolidates the $N\times N$ covariance matrices into $N^{2}$-dimensional feature vectors for each time step within the sequence.
2.11.3 Bidirectional LSTM (BiLSTM) Layer
Following the CNN, the flattened transformed sequences are fed into a layer configured with $u$ layers and $h_{d}$ hidden dimension. To understand the BiLSTM, it is essential first to grasp the fundamentals of a standard LSTM network (Hochreiter,, 1997). Appendix A.1 provides more detailed information on how each LSTM cell works.
The BiLSTM enhances the standard LSTM by adding another layer that processes the input sequence in the reverse order. This bidirectional approach allows the model to have both forward and backward information about the sequence, making it more effective in capturing the context from past and future data points. Each sequence goes through $u$ layers, i.e. the output of layer one serves as input for layer 2.
In our model, the BiLSTM first transforms the input sequence $\mathbf{X}_{\text{flattened}}$ of dimension $L\times N^{2}$ into an output sequence of dimension $L\times 2h_{d}$. The BiLSTM uses learnable parameters to map the $N^{2}$ features at each time step into a lower-dimensional hidden representation of size $h_{d}$. This transformation is achieved through linear transformations, where weights and biases are optimised during training. Then, the hidden states from both forward and backward pass across all layers, and time steps are concatenated to form a comprehensive hidden state matrix. This matrix captures information from all directions and spans the entire sequence of inputs. Each BiLSTM layer can be expressed as:
$\mathbf{H}^{o}=\begin{bmatrix}\mathbf{H}^{(1)},\mathbf{H}^{(2)},\dots,\mathbf{H}^{(h_{d})}\\ \end{bmatrix}\in\mathbb{R}^{L\times 2h_{d}}$ (31)
where $o$ represents the BiLSTM layer, for $o=1,\dots,u$, and each $\mathbf{H}^{(l)}$ for $l=1,2,\ldots,h_{d}$ is defined as:
$\mathbf{H}^{(l)}=\begin{bmatrix}\overrightarrow{\mathbf{h}}^{(l)}_{t-L}&\overleftarrow{\mathbf{h}}^{(l)}_{t-L}\\ \overrightarrow{\mathbf{h}}^{(l)}_{t-L+1}&\overleftarrow{\mathbf{h}}^{(l)}_{t-L+1}\\ \vdots&\vdots\\ \overrightarrow{\mathbf{h}}^{(l)}_{t}&\overleftarrow{\mathbf{h}}^{(l)}_{t}\\ \end{bmatrix}\in\mathbb{R}^{L\times 2}$ (32)
The hidden dimension for each LSTM direction, denoted as $h_{d}$, results in a concatenated hidden state dimension of $2h_{d}$ at each time step.
The traditional BiLSTM outputs the last time step of the last layer. However, in our model, we use a similar approach to Liu and Guo, (2019) methodology, meaning all time steps from the last layer ($\mathbf{H}^{u}$) are used as an input to our following steps. To mitigate overfitting and enhance the model’s generalisation capabilities, a Dropout Layer is applied to the BiLSTM output. This operation randomly zeroes a fraction of the elements in $\mathbf{H}^{u}$ with a probability of 20%, effectively regularising the model by preventing reliance on specific neurons during training. During our testing phase, this feature will be skipped.
2.11.4 Multi-Head Self-Attention Layer
Our next step is to apply a multi-head attention mechanism (Vaswani et al.,, 2017) to each hidden state of the last layer. The multi-head attention mechanism to refine the temporal dependencies captured in the sequence. For $h_{e}$ attention heads, the input $\mathbf{H}^{u}$ is used to compute the queries ($\mathbf{Q}$), keys ($\mathbf{K}$), and values ($\mathbf{V}$): $\mathbf{Q}=\mathbf{H}^{u}\textbf{W}_{q}$, $\mathbf{K}=\mathbf{H}^{u}\textbf{W}_{k}$, and $\mathbf{V}=\mathbf{H}^{u}\textbf{W}_{v}$, where $\textbf{W}_{q},\textbf{W}_{k},\textbf{W}_{v}\in\mathbb{R}^{2h_{d}\times 2h_{d}}$ are learnable parameters. The attention weights for each head are computed as follows:
$\textbf{Attention}(\mathbf{Q}_{\text{head}},\mathbf{K}_{\text{head}},\mathbf{V}_{\text{head}})=\text{softmax}\left(\frac{\mathbf{Q}_{\text{head}}\mathbf{K}_{\text{head}}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{\text{head}}\in\mathbb{R}^{L\times d_{k}}$ (33)
where $d_{k}=\frac{2h_{d}}{h_{e}}$ scales the dot product to stabilise gradients and splits the features. Each head processes its respective portion of the features, which are concatenated to produce $\mathbf{A}_{concat}\in\mathbb{R}^{L\times 2h_{d}}$. The concatenated output is projected back to $\mathbf{A}=\mathbf{A}_{concat}\textbf{W}_{c}\in\mathbb{R}^{L\times 2h_{d}}$ using a linear transformation.
The attention output is mean-pooled along the sequence dimension to condense the temporal information into a fixed-length representation:
$\mathbf{a}_{\text{mean}}=\frac{1}{L}\sum_{t=1}^{L}\mathbf{A}_{t}\quad\in\mathbb{R}^{1\times 2h_{d}}$ (34)
where $\mathbf{A}_{t}$ is the $t$-th row of $\mathbf{A}$. This averaged context vector $\mathbf{a}_{\text{mean}}$ encapsulates the overall information from the entire sequence. We then pass $\mathbf{a}_{\text{mean}}$ through a fully connected layer to map it back to the covariance matrix space:
This averaged context vector $\mathbf{a}_{\text{mean}}$ encapsulates the overall information from the entire sequence. We then pass $\mathbf{a}_{\text{mean}}$ through a fully connected layer to map it back to the covariance matrix space:
$\mathbf{y}=\mathbf{a}_{\text{mean}}\mathbf{W}_{\text{fc}}+\mathbf{b}_{\text{fc}}\quad\in\mathbb{R}^{1\times N^{2}}$ (35)
where $\mathbf{W}_{\text{fc}}$ and $\mathbf{b}_{\text{fc}}$ are receptively the weight matrix and bias vector of the fully connected layer.
Because our pipeline stacks 3D Convolution, BiLSTM, and attention in succession, the final latent representation becomes highly abstract, making interpretability at each layer difficult. Consequently, we do not attempt to interpret the hidden states or attention weights in a domain sense.
2.11.5 Enforcing Symmetry and Positive Semi-Definiteness
The output vector $\mathbf{y}$ is reshaped into a square matrix ($\mathbf{Y}_{\text{reshaped}}\in\mathbb{R}^{N\times N}$) to form the predicted covariance matrix.
The output matrix $\mathbf{Y}_{\text{reshaped}}$ may not necessarily be symmetric, a key property of covariance matrices. To enforce symmetry, we average the matrix with its transpose:
$\mathbf{Y}_{\text{sym}}=\frac{\mathbf{Y}_{\text{reshaped}}+\mathbf{Y}_{\text{reshaped}}^{\top}}{2}$ (36)
Then, we rescale the normalised output back to the original scale of the data, using the reverse process of Section 2.11.1 resulting in $\mathbf{Y}_{\text{sym}}^{unscaled}$.
Symmetry alone does not guarantee that $\mathbf{Y}_{\text{sym}}^{unscaled}$ is positive semi-definite, which is another requisite property of covariance matrices. To ensure positive semi-definiteness, we perform eigenvalue decomposition:
$\mathbf{Y}_{\text{sym}}^{unscaled}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$ (37)
We then modify $\mathbf{\Lambda}$ by setting all negative eigenvalues to zero:
$\mathbf{\Lambda}^{\prime}=\text{diag}\left(\max(\lambda_{i},0)\right)$ (38)
The positive semi-definite matrix is reconstructed as:
$\mathbf{Y}_{\text{psd}}=\mathbf{U}\mathbf{\Lambda}^{\prime}\mathbf{U}^{\top}$ (39)
2.11.6 Shrinkage
While DL-based forecasts can learn intricate patterns, historical covariance may still convey valuable stability and continuity, particularly when the model is tested on data different from training. Therefore, our model’s final step is incorporating a linear shrinkage from historical values in our predictions.
$\textbf{CAB}=\mathbf{Y}_{\text{psd}}\dot{\phi}+\mathbf{\Sigma}_{t-F:t}\dot{(}1-\phi)$ (40)
where CAB is the final $N\times N$ matrix output, and $\phi$ is the shrinkage factor.
2.12 Evaluation metrics
We will use performance metrics consistent with previous literature (Symitsi et al.,, 2018; Bollerslev et al.,, 2018; Zhang et al.,, 2024) to check the models’ performance. We consider the following loss functions to measure the average distance between predicted covariances and realised covariances matrices for the model comparisons:
$\mathcal{L}_{t}^{E}=\sqrt{\operatorname{vec}_{Tri}(\mathbf{\Sigma}_{t}-\hat{\mathbf{\Sigma}}_{t})^{\top}\operatorname{vec}_{Tri}\left(\mathbf{\Sigma}_{t}-\mathbf{\hat{\Sigma}}_{t}\right)}$ (41)
$\mathcal{L}_{t}^{F}=\sqrt{\operatorname{Tr}\left[\left(\bm{\Sigma}_{t}-\bm{\hat{\Sigma}}_{t}\right)^{\top}\left(\bm{\Sigma}_{t}-\bm{\hat{\Sigma}}_{t}\right)\right]}$ (42)
where $\mathcal{L}_{t}^{E}$ represents the Euclidean distance between the $N(N+1)/2$ dimensional vectorised version of the upper triangular of forecast covariances and ex-post realised covariances. $\mathcal{L}_{t}^{F}$ is the Frobenius distance between the two matrices, and $\operatorname{Tr}$ is the trace of a square matrix. All these functions measure losses. Therefore, lower values are preferred.
3 Empirical Results
We obtain our sample data from the Refinitiv Database, covering daily adjusted closing prices of 14 major ETFs that reflect equity and bond markets across diverse global sectors. These ETFs were selected to construct a multi-asset environment that includes equity sectors (e.g. technology, financials, industrials, consumer staples) and fixed-income exposures. Table 1 details the ETF’s information. To compute excess returns, we also retrieve from the Federal Reserve Economic Data the 1-month U.S. Treasury yield, treating it as a daily risk-free rate. Our sample runs from 1 January 2017 to 31 December 2023, totalling 1,760 trading days.
Table 1: List of ETFs included in our research.
Name Ticker ISIN Asset Class
iShares Global Tech ETF IXN US4642872919 Equity
iShares Global Financials ETF IXG US4642873339 Equity
iShares Global Consumer Discretionary ETF RXI US4642887453 Equity
iShares Global Industrials ETF EXI US4642887297 Equity
iShares Global Healthcare ETF IXJ US4642873255 Equity
iShares Global Consumer Staples ETF KXI US4642887370 Equity
iShares U.S. Real Estate ETF IYR US4642877397 Real Estate
iShares International Developed Real Estate ETF IFGL US4642884898 Real Estate
iShares Global Materials ETF MXI US4642886950 Equity
iShares Global Energy ETF IXC US4642873412 Equity
iShares Global Comm Services ETF IXP US4642872752 Equity
iShares Global Utilities ETF JXI US4642887115 Equity
iShares Core U.S. Aggregate Bond ETF AGG US4642872265 Fixed Income
iShares Core International Aggregate Bond ETF IAGG US46435G6724 Fixed Income
We split our dataset into a training period from 1 January 2017 to 31 December 2020 and a testing period covering 1 January 2021 to 31 December 2023. The testing sample thus comprises 753 trading days, encompassing a wide range of economic and market conditions. We identify three distinct market regimes within this testing window to examine whether our model responds robustly to varying interest-rate and equity-market environments. We designate the interval between 1 January 2021 and 2 January 2022 as the First Bull Period (Bull-1), typified by low yields on the risk-free asset and a generally stable equity market. We label the interval from 3 January 2022 to 12 June 2022 as the Bear Period (Bear), characterised by a steady rise in the risk-free rate and a drop of 20% or more in the S&P 500 from its last peak. Lastly, we classify the interval from 13 June 2022 to 31 December 2023 as the Second Bull Period (Bull-2), featuring persistently high interest rates but a renewed upswing in equity prices. This subdivision ensures that our out-of-sample evaluations capture both low- and high-rate market phases and a bear-market regime, offering a comprehensive test of the model’s ability to adapt to evolving financial landscapes.
During the training period, we minimise the Frobenius distance (Eq. 42) to select hyperparameters. We use Optuna (Akiba et al.,, 2019) for hyperparameter selection. Appendix A.2 shows the grid search table. Specifically, we split the training dataset into a model-fitting subset (80%) and a validation subset (20%) to identify optimal settings under a forecast horizon ($F$) of 20 days. We adopt the Adam optimiser over 100 epochs, using a batch size of 128 and a learning rate of $10^{-4}$. For the neural-network-based $CAB$ model, we use a lookback window ($L$) of 100 days. We set the 3D convolution kernel size ($ks$) to five; the BiLSTM has 128 hidden dimensions ($hd$) across seven stacked layers ($u$), and the multi-head attention uses 16 heads ($he$). Finally, we apply a shrinkage factor of 0.8 ($\phi$) when combining the DL output with the most recent historical covariance.
We conduct a rolling estimation procedure to simulate live conditions. Each new trading day $t$ in the test set triggers an online update of our DL model, incorporating all data up through day $t$. This design ensures that the forecasts mimic a real-world trading environment, where only up-to-date historical data is available at each decision point.
Table 2 reports the average Euclidean ($\mathcal{L}^{E}$) and Frobenius ($\mathcal{L}^{F}$) distances between forecasted and realised covariance matrices for all models over the entire testing period (2021–2023) [^4] Bold and underlined numbers represent the lowest and the second lowest values, respectively.. GARCH models ($CCC$ and $DCC$) and $EWMA$ perform relatively well as they adapt faster to structural changes in the covariance matrices. However, our model ($CAB$) yields the lowest errors in both metrics, outperforming classical benchmarks. Dimensional reduction models perform inadequately as we only use 14 assets.
Table 2: Average Euclidean ($\mathcal{L}^{E}$) and Frobenius ($\mathcal{L}^{F}$) distances (×$\times 10^{5}$) for whole the testing period (1 January 2021 to 31 December 2023).
$\mathcal{L}^{E}$ $\mathcal{L}^{F}$
$NA$ 53.526 82.020
$NA^{F}$ 52.713 80.987
$EWMA$ 49.111 75.349
$LW$ 49.772 76.514
$LW^{F}$ 68.136 103.363
$CCC$ 48.568 75.198
$DCC$ 48.268 74.716
$DCC^{NL}$ 51.621 79.079
$PCA$ 54.348 83.154
$RMT$ 54.145 82.915
$CAB$ 38.312 58.923
To solidify our findings, we analysed the 10 models with 753 paired samples for the Euclidean and Frobenius distances following Demsar, (2006) frequentist methodology. The significance level is $\alpha=0.050$ for all tests here forward. We rejected the null hypothesis that the population is normal for all the populations ($p<0.01$ for all models in both distances). Therefore, we assume that not all populations are normal.
Because we have more than two populations and the populations are not normal, we use the non-parametric Friedman test as an omnibus test to determine if there are any significant differences between the median values of the populations. We use the post-hoc Nemenyi test to infer which differences are significant. Differences between populations are significant if the difference in the mean rank is greater than the critical distance CD=0.494 of the Nemenyi test.
We reject the Friedman test’s null hypothesis, which states that there is no difference in the central tendency of the populations. Therefore, we assume that there is a statistically significant difference between the central tendencies of the populations. Figure 1 shows the post-hoc Nemenyi test critical distance between groups. Our model reports statistically better results than the benchmark models.
(a) Upper Triangular Euclidean distance
(b) Frobenius distance
Figure 1: Critical distance diagrams for the post-hoc Nemenyi test.
Table 3 presents the average distance metrics for the testing period across three market regimes: Bull-1, Bear, and Bull-2. Overall, the $CAB$ model consistently outperforms all benchmarks.
Table 3: Average Euclidean ($\mathcal{L}^{E}$) and Frobenius ($\mathcal{L}^{F}$) distances (×$\times 10^{5}$) for different testing periods: Bull-1 (1 January 2021 to 2 January 2022), Bear (3 January 2022 to 12 June 2022), and Bull-2 (13 June 2022 to 31 December 2023).
Bull-1 Bear Bull-2
$\mathcal{L}^{E}$ $\mathcal{L}^{F}$ $\mathcal{L}^{E}$ $\mathcal{L}^{F}$ $\mathcal{L}^{E}$ $\mathcal{L}^{F}$
$NA$ 36.937 56.621 89.891 137.868 53.709 82.252
$NA^{F}$ 35.680 54.992 98.002 150.780 50.586 77.545
$EWMA$ 31.754 48.973 82.799 126.614 50.574 77.549
$LW$ 34.126 52.599 90.441 138.173 48.089 74.088
$LW^{F}$ 63.288 94.759 93.189 143.636 63.989 97.222
$CCC$ 31.688 49.784 88.106 134.902 48.017 74.316
$DCC$ 33.516 52.166 84.745 130.345 47.227 73.161
$DCC^{NL}$ 36.779 56.449 87.650 134.167 50.767 77.734
$PCA$ 37.424 57.281 90.720 139.116 54.747 83.660
$RMT$ 37.246 57.083 90.638 138.821 54.493 83.411
$CAB$ 24.276 37.643 66.381 102.252 39.253 60.126
While all models experience a decrease in accuracy during the turbulent Bear regime, the $CAB$ approach exhibits notable robustness. In particular, during the Bull-2 period, characterised by elevated interest rates and dynamic shifts in cross-asset relationships, integrating DL techniques with classical shrinkage enables $CAB$ to adapt effectively to changing market conditions, thereby maintaining lower distance metrics.
Supplementary Nemenyi post-hoc tests confirm that the performance improvements observed with $CAB$ are statistically significant. These tests provide further evidence that the model’s superior performance is not merely a result of sample-specific idiosyncrasies but reflects a robust ability to handle varying market regimes.
Moreover, these findings contribute to the literature on hybrid forecasting methods by demonstrating that blending modern DL with traditional econometric techniques can offer substantial gains in predictive accuracy, especially in periods of market stress. This robust performance under diverse macroeconomic conditions highlights the practical relevance of the $CAB$ approach for risk management and strategic asset allocation.
4 Robustness analysis
4.1 Impact of Using Raw Returns vs. Excess Returns
One potential criticism of our model is that it may be overly tailored to excess returns, even though some classic optimisation frameworks (Markowitz,, 1952) often use raw returns. We re-estimate our model without subtracting the risk-free rate to address this concern. All other model parameters remain identical to those described in Section 2, with the sole difference being that $\mathbf{r}_{t}$ now represents the $N$-dimensional row vector of raw daily returns for $N$ assets at time $t$.
Table 4 reports the performance of the various models under these raw returns assumptions. Our approach ($CAB$) continues outperforming all benchmarks across the entire sample and within individual market regimes, with statistically significant results based on the same post-hoc Nemenyi criteria used earlier. This finding directly informs our second research question about whether raw vs. excess returns lead to materially different covariance forecasts. When comparing Table 4 (raw returns) to Tables 2 and 3 (excess returns), the numerical magnitudes change slightly, but the rank order remains the same: $CAB$ still produces the most accurate forecasts overall.
Our findings indicate that the advantage of our framework is not contingent on the choice between excess or raw returns, reaffirming its robustness to the presence or absence of the risk-free rate. Moreover, this consistency suggests that subtracting the risk-free rate over the medium-term horizon does not significantly alter the core structure of asset co-movements. Hence, even though economic theory posits that excess returns better capture risk premiums, our empirical results show that the forecasting edge of $CAB$ holds across both raw and excess return specifications. In other words, our method retains its superior performance even when modelling raw returns, suggesting that its effectiveness does not hinge on subtracting the risk-free rate.
Table 4: Average Euclidean ($\mathcal{L}^{E}$) and Frobenius ($\mathcal{L}^{F}$) distances (×$\times 10^{5}$) using raw returns for different testing periods: Overall (1 January 2021 to 31 December 2023), Bull-1 (1 January 2021 to 2 January 2022), Bear (3 January 2022 to 12 June 2022), and Bull-2 (13 June 2022 to 31 December 2023).
Overall Bull-1 Bear Bull-2
$\mathcal{L}^{E}$ $\mathcal{L}^{F}$ $\mathcal{L}^{E}$ $\mathcal{L}^{F}$ $\mathcal{L}^{E}$ $\mathcal{L}^{F}$ $\mathcal{L}^{E}$ $\mathcal{L}^{F}$
$NA$ 53.525 82.020 36.936 56.620 89.895 137.875 53.707 82.249
$NA^{F}$ 52.486 80.668 34.913 53.883 98.623 151.717 50.462 77.372
$EWMA$ 49.174 75.441 31.753 48.972 82.798 126.613 50.697 77.730
$LW$ 49.771 76.512 34.125 52.598 90.442 138.174 48.087 74.085
$LW^{F}$ 68.130 103.353 63.273 94.736 93.190 143.636 63.985 97.217
$CCC$ 48.508 75.106 31.599 49.634 88.139 135.038 47.948 74.196
$DCC$ 48.232 74.661 33.445 52.045 84.884 130.630 47.161 73.050
$DCC^{NL}$ 51.604 79.051 36.691 56.305 87.856 134.535 50.731 77.664
$PCA$ 54.348 83.153 37.423 57.280 90.724 139.123 54.745 83.657
$RMT$ 54.144 82.914 37.245 57.081 90.643 138.829 54.491 83.408
$CAB$ 36.663 56.518 22.807 35.673 62.036 96.406 38.273 58.441
4.2 Varying Forecast Horizons
We also examine our model’s behaviour as the forecast horizon ($F$) varies. In the baseline analysis, $F=20$ days are medium-term, but many institutional investors face different time horizons for risk management or asset allocation decisions. In line with our first research question regarding underexplored medium-term horizons, we extend $F$ from 10 to 250 days, thus covering short, medium, and long horizons. To explore model performance across this range of settings, we evaluate both excess (Table LABEL:tab:diffhorizonsexc) and raw returns (Table LABEL:tab:diffhorizonsraw).
Table LABEL:tab:diffhorizonsexc shows that our model consistently achieves the lowest average distance metrics across most forecast horizons. $CAB$ remains among the top performers for shorter horizons ($F=10$). However, it does not outperform GARCH-based models ($CCC$ and $DCC$) and $EWMA$. These differences are not statistically different in the overall period, Bull-1 and Bull-2. During Bull-2, our model outperformed the benchmarks, and the results were statistically significant based on the post-hoc Nemenyi test.
For mid-term forecasting ($F=40,60,90,120,180$) for both distances, the differences are statistically significant in all periods. Thus, our model consistently outperforms the benchmarks. Please note that GARCH models tend to underperform as the forecast horizon increases, while full data models ($NA^{F}$ and $LW^{F}$) tend to diminish forecasting errors. Please note, while the $DCC^{NL}$ was more recently designed for high dimensional matrices, it provides more stable forecasts based on eigenvalues than the classical $DCC$.
At the longest horizon ($F=250$), $CAB$ again excels in most periods. Besides showing a large difference between distances during the Bear period and Bull-1, the post-hoc Nemenyi test shows no statistical difference between our model and $NA^{F}$ and $LW^{F}$, respectively. These results suggest that more complex models capture slower-moving cross-asset dynamics effectively, while classical estimators remain competitive in certain regime-specific conditions. As anticipated, naive and shrinkage methods perform relatively well at the longer horizon. These patterns reinforce our first research question’s premise that medium-term horizons can benefit from sophisticated models and highlight instances where simpler approaches hold their own.
Turning to Table LABEL:tab:diffhorizonsraw, we link our findings to the second research question: whether using raw or excess returns makes a systematic difference. Results show that the top-performing models remain consistent across all forecast horizons.
Once again, $CAB$ tends to yield the strongest overall performance. Although it does not dominate at shorter horizons, it consistently ranks near the top, showing no statistical difference between our model and the top performers. All other differences are statistically significant. These results mirror the trends seen under excess returns, indicating that the performance edge of $CAB$ does not hinge on subtracting the risk-free rate. Therefore, our DL approach remains robust even if practitioners opt for raw returns, common in certain classical optimisation frameworks. Our model performs better than the benchmarks for longer horizons ($F=250$). However, the results are not statistically different from the second-performing models ($NA^{F}$ and $LW^{F}$) for the Bull-1 and Bear periods, respectively.
These findings indicate that DL-based models can adapt effectively across different forecast windows. Across both dimensions, $CAB$ shows a stable advantage, with only minor variations in extreme horizons.
5 Economic Value in Portfolio Management
Next, we investigate the economic importance of accurate covariance forecasts in a practical asset allocation framework. Specifically, we consider an investor allocating wealth among the ETFs specified in section 3, subject to a no-short-selling constraint. Such a constraint is common among regulated institutional investors (e.g., pension funds, mutual funds) that may face limits on leverage or short positions. We adopt daily, weekly, and monthly rebalancing intervals (Symitsi et al.,, 2018).
At each rebalancing date, the investor solves the GMV problem:
$\min_{\bm{w}_{t}}\quad\mathbf{w}_{t}^{\top}\,\widehat{\mathbf{\Sigma}}_{t}\,\mathbf{w}_{t}\quad\text{subject to}\quad\mathbf{1}^{\top}\mathbf{w}_{t}=1,\quad w_{i,t}\geq 0,$ (43)
where $\mathbf{w}_{t}$ is an $N\times 1$ vector of GMV portfolio weights, $\widehat{\mathbf{\Sigma}}_{t}$ is the $N\times N$ forecast covariance from each model, and $\mathbf{1}$ is an $N\times 1$ vector of ones. Excluding expected returns from the optimisation allows us to isolate the impact of covariance forecasts (DeMiguel et al.,, 2009; Kourtis et al.,, 2012).
Using a rolling estimation window, we generate 11 portfolio strategies (one for each forecast model) plus an equally weighted (1/$N$) benchmark. After estimating $\widehat{\mathbf{\Sigma}}_{t}$, we compute the GMV weights and hold them until the next rebalancing date. We then record the ex-post average portfolio return for each model ($m$) as:
$\mathbf{r}^{(m)}_{t+1}=\frac{1}{P}\sum_{t=1}^{P}\mathbf{w}^{(m)}_{t}{}^{\top}\,\mathbf{r}_{t+1},$ (44)
where $\mathbf{r}_{t}$ is an $N\times 1$ vector of realised asset returns, $P$ is the total out-of-sample length, and $\mathbf{w}^{(m)}_{t}$ are the portfolio weights at the start of period $t$ for model $m$.
We focus on two main metrics to compare forecast models: out-of-sample portfolio variance and turnover. The variance of daily returns by a portfolio constructed by model $m$ is:
$\sigma_{m}^{2}=\frac{1}{P}\sum_{t=1}^{P}\bigl{(}r^{(m)}_{t}-\bar{r}^{(m)}\bigr{)}^{2}$ (45)
where $r^{(m)}_{t}$ is the ex-post return of model $m$ in period $t$, $\bar{r}^{(m)}$ and $P$ is its mean and number of observations over the out-of-sample window. Meanwhile, the average turnover reflects how often the portfolio adjusts its weights:
$\text{TO}_{m}=\frac{1}{P-1}\sum_{t=1}^{P-1}\bigl{\lVert}\mathbf{w}_{t+1}^{(m)}-\mathbf{w}_{t+}^{(m)}\bigr{\rVert}_{1},$ (46)
where $\mathbf{w}_{t+}^{(m)}$ are the portfolio weights of model $m$ immediately before rebalancing at $t+1$, and $\lVert\cdot\rVert_{1}$ is the 1-norm. A higher turnover implies more frequent portfolio adjustments, potentially incurring greater trading costs (Kourtis,, 2014).
We also include an equally weighted (1/$N$) portfolio, often cited for its simplicity and historically strong performance (DeMiguel et al.,, 2009). Although $1/N$ avoids estimation risk and typically maintains low turnover, it cannot adapt to shifting market correlations.
From Table 5, it is evident that the covariance forecasts provided by the $CAB$ model strike a competitive balance between risk reduction and turnover across all rebalancing horizons. Across daily, weekly, and monthly rebalancing horizons, $CAB$ slightly trails the top daily performer, approaches the leading weekly method, and matches or surpasses most other approaches monthly, thus consistently ranking among the best-performing models. This robust performance does not come at the cost of excessive turnover, as CAB’s turnover measures remain moderate compared to methods that may obtain marginally lower variances only at the expense of frequent portfolio rebalancing (e.g., $NA$, $DCC$, $DCC^{NL}$).
Table 5: performance of the global minimum variance portfolios constructed using the covariance forecasts from the 11 models under consideration plus the equal-weighted portfolio. The portfolios are compared based on their annualised out-of-sample standard deviation ($\sigma_{m}$) and average out-of-sample turnover ($\text{TO}_{m}$), respectively. Results for three different rebalancing frequencies are presented.
Daily Weekly Monthly
$\sigma_{m}^{2}$ $TO_{m}$ $\sigma_{m}^{2}$ $TO_{m}$ $\sigma_{m}^{2}$ $TO_{m}$
$1/N$ 0.018132 0.005733 0.018115 0.002567 0.018060 0.001282
$NA$ 0.001991 0.125568 0.002527 0.055395 0.002631 0.025707
$NA^{F}$ 0.002451 0.001445 0.002453 0.000653 0.002461 0.000369
$EWMA$ 0.002475 0.079388 0.002482 0.037928 0.002554 0.016088
$LW$ 0.002655 0.098020 0.003373 0.047089 0.003598 0.022957
$LW^{F}$ 0.002774 0.001816 0.002777 0.000956 0.002784 0.000620
$CCC$ 0.002281 0.126519 0.002580 0.049839 0.002713 0.014816
$DCC$ 0.002075 0.116446 0.002388 0.051710 0.002524 0.017159
$DCC^{NL}$ 0.002048 0.129453 0.002448 0.056173 0.002618 0.023470
$PCA$ 0.002680 0.197322 0.002722 0.072903 0.002657 0.023108
$RMT$ 0.002647 0.125512 0.002689 0.055954 0.002733 0.024767
$CAB$ 0.002293 0.078799 0.002428 0.048383 0.002579 0.023807
When compared to the $1/N$ benchmark, please note that all models present a statistically significant lower annual variance, based on the one-side F-Test of equalities of variance.
Examining the same metrics across distinct market regimes (Table LABEL:tab:portfoliovariancesmr) reveals that $CAB$ remains a robust contender under varying conditions: in Bull-1, it consistently competes for the lowest variance, even attaining the best performance in weekly and monthly rebalancing; in the Bear regime, $CAB$ ranks near the top in daily and weekly horizons, balancing lower turnover against marginally higher volatility than a few dynamic methods; and while Bull-2 sees other estimators occasionally surpass $CAB$, it still demonstrates stable risk control with moderate trading activity, indicating that its effectiveness persists across both tranquil and turbulent market environments.
Comparisons to the $1/N$ benchmark further highlight the importance of adaptivity. While $1/N$ offers minimal turnover, it endures noticeably higher variance in all market conditions, reflecting its inability to exploit up-to-date correlation information. Indeed, the difference in annualised variances between $CAB$ and $1/N$ can reach 80-90 basis points in certain scenarios, a gap that can substantially affect risk-adjusted returns and capital preservation.
Overall, $CAB$ tends to exhibit moderate turnover, balancing the need to adjust to changing correlations without incurring unnecessary trades. This stability benefits investors who must limit transaction costs or adhere to regulatory constraints on turnover. These results underscore that advanced covariance estimation can yield economically meaningful improvements in risk control, particularly over the medium-term horizons favoured by many institutional investors. The $CAB$ model stands out by combining strong variance reduction with stable rebalancing behaviour, balancing responsiveness to market changes and turnover efficiency.
6 Conclusion
We propose a novel DL framework for medium-term covariance forecasting in multi-asset portfolios, combining 3D convolutions, bidirectional LSTMs, and multi-head attention. Empirical tests on various ETFs from 2017 to 2023 reveal that this model, enhanced by a final shrinkage step, consistently outperforms classical benchmarks across different forecast horizons and market regimes.
Notably, its performance advantage remains robust whether one subtracts a risk-free rate from returns, suggesting broad applicability in diverse portfolio management practices.
In portfolio experiments, the proposed method enables global minimum-variance (GMV) strategies to achieve lower out-of-sample volatility with moderate turnover, underlining the tangible economic value of improved covariance estimation. By bridging cutting-edge DL techniques with established financial principles, our work highlights the promise of sophisticated spatiotemporal modelling for risk management and allocation decisions, especially at horizons where structural shifts and evolving correlations pose unique forecasting challenges.
Still, several research avenues warrant further exploration. First, extending the network architecture to accommodate skewed or heavy-tailed distributions could better capture tail risk in times of market stress. Second, explicitly incorporating transaction costs or liquidity constraints within the optimisation process might yield more realistic and implementable trading strategies. Finally, evaluating the model in larger cross-sectional settings or applied to alternative asset classes (e.g., commodities, crypto assets) would illuminate its scalability and robustness.
We hope these findings encourage the finance community to explore and adapt advanced neural architectures for medium-term and long-horizon applications, strengthening the link between machine learning innovation and effective real-world asset allocation.
Appendix A Appendix
A.1 LSTM cell detailed explanation
An LSTM (Hochreiter,, 1997) cell consists of three main gates: the forget gate ($\textbf{{f}}_{t}$), the input gate ($\textbf{{i}}_{t}$), and the output gate ($\textbf{{o}}_{t}$). They are respectively calculated as:
$\textbf{{f}}_{t}=\sigma(\mathbf{W}_{f}.[\mathbf{h}_{(t-1)},\mathbf{x}_{t}\ ]+\mathbf{b}_{f})$ (A.1)
$\textbf{{i}}_{t}=\sigma(\mathbf{W}_{i}.[\mathbf{h}_{(t-1)},\mathbf{x}_{t}\ ]+\mathbf{b}_{i})$ (A.2)
$\textbf{{o}}_{t}=\sigma(\mathbf{W}_{o}.[\mathbf{h}_{(t-1)},\mathbf{x}_{t}\ ]+\mathbf{b}_{o})$ (A.3)
where $\mathbf{x}_{t}$ is the vector input at time $t$, $\mathbf{h}_{(t-1)}$ is the hidden vector state, $\mathbf{W}_{f}$, $\mathbf{W}_{i}$ and $\mathbf{W}_{c}$ are weight matrices, and $\mathbf{b}_{f}$, $\mathbf{b}_{i}$ and $\mathbf{b}_{o}$ are bias vectors. The sigmoid function $\sigma$ ensures that the output values are between 0 and 1, representing how much each component should be forgotten.
The forget gate decides what information should be discarded from the cell state. Next, the input gate determines what new information should be added to the cell state. This process involves two steps: calculating the input gate Eq. (A.2) and creating new candidate values:
$\mathbf{\widetilde{C}}_{t}=\tanh\left(\mathbf{W}_{c}\cdot\left[\mathbf{h}_{t-1},\mathbf{x}_{t}\right]+\mathbf{b}_{c}\right)$ (A.4)
where $\mathbf{W}_{c}$ is the weight matrix, and $\mathbf{b}_{c}$ is a bias vector. The input gate modulates the extent to which new information is added to the cell state. The cell state update combines the forget gate and input gate operations to update the cell state:
$\mathbf{C}_{t}=\textbf{{f}}_{t}\odot\mathbf{C}_{(}t-1)+\textbf{{i}}_{t}\odot\mathbf{\widetilde{C}}_{t}$ (A.5)
This equation ensures that the cell state retains essential information over long periods.
Finally, the output gate controls what information should be outputted from the cell. This involves two steps: calculating the output gate Eq. (A.3) and determining the hidden state:
$\mathbf{h}_{t}=\textbf{{o}}_{t}\odot\tanh{\left(\mathbf{C}_{t}\right)}$ (A.6)
A.2 Hyperparameter Grid Search
Table A1: Hyperparameter Grid Search table.
Hyperparameter Range
Optimiser Adam, SGD
Batch size 32, 64, 128, 256
Learning rate $10^{-3}$, $10^{-4}$, $10^{-5}$
Lookback window ($L$) 20, 40, 60, 80, 100, 120, 250
3D convolution kernel size ($ks$) 3, 5, 7
BiLSTM hidden dimensions ($hd$) 32, 64, 128, 256
BiLSTM stacked layers ($u$) 3, 4, 5, 6, 7
Number of heads ($he$)a 2, 4, 6, 8, 16, 32
Shrinkage factor ($\phi$) 0.00, 0.20, 0.40, 0.60, 0.80, 1.00
A.3 Extended Data Tables
Table A2: Average Euclidean ($\mathcal{L}^{E}$) and Frobenius ($\mathcal{L}^{F}$) distances (×$\times 10^{5}$) using excess returns with different horizons ($F$) for different testing periods: Overall (1 January 2021 to 31 December 2023), Bull-1 (1 January 2021 to 2 January 2022), Bear (3 January 2022 to 12 June 2022), and Bull-2 (13 June 2022 to 31 December 2023).
Overall Bull-1 Bear Bull-2
$\mathcal{L}^{E}$ $\mathcal{L}^{F}$ $\mathcal{L}^{E}$ $\mathcal{L}^{F}$ $\mathcal{L}^{E}$ $\mathcal{L}^{F}$ $\mathcal{L}^{E}$ $\mathcal{L}^{F}$
$F=10$ $NA$ 70.309 107.739 54.194 82.600 107.447 165.092 69.960 107.363
$NA^{F}$ 62.260 95.586 46.956 71.921 106.652 163.471 59.271 91.185
$EWMA$ 58.851 90.253 42.125 64.612 94.704 145.253 59.272 90.889
$LW$ 63.548 97.681 47.376 72.757 104.804 160.926 62.036 95.449
$LW^{F}$ 75.861 115.370 69.194 104.027 105.832 162.356 71.461 109.054
$CCC$ 58.111 89.824 42.451 65.818 95.819 147.003 57.301 88.763
$DCC$ 62.210 95.018 48.557 73.797 96.913 147.989 60.972 93.374
$DCC^{NL}$ 65.361 99.154 51.461 77.640 100.48 152.649 64.160 97.548
$PCA$ 71.266 109.164 54.984 83.752 110.811 170.301 70.323 107.864
$RMT$ 71.064 108.951 54.819 83.588 110.623 170.011 70.094 107.641
$CAB$ 56.954 87.317 44.134 67.401 94.689 145.287 54.291 83.369
$F=40$ $NA$ 45.872 70.272 24.755 38.850 72.135 108.838 51.939 79.449
$NA^{F}$ 47.436 72.906 31.056 47.811 90.540 139.813 45.522 69.719
$EWMA$ 44.734 68.528 26.175 40.546 74.009 112.405 48.262 73.923
$LW$ 43.441 66.728 23.834 37.275 76.162 114.702 46.645 71.886
$LW^{F}$ 63.848 96.730 62.374 93.125 78.628 122.264 60.501 91.635
$CCC$ 43.434 67.183 27.608 43.493 75.190 114.891 44.465 68.675
$DCC$ 42.622 66.124 27.194 42.996 73.983 113.348 43.509 67.391
$DCC^{NL}$ 43.764 67.604 28.135 44.201 75.173 114.919 44.767 69.024
$PCA$ 46.390 71.007 24.934 39.238 72.294 109.567 52.781 80.410
$RMT$ 46.177 70.771 24.796 39.028 72.411 109.269 52.424 80.176
$CAB$ 22.851 35.411 13.091 20.268 42.372 65.754 23.505 36.408
$F=60$ $NA$ 44.053 68.114 26.115 42.593 72.110 109.681 47.531 72.584
$NA^{F}$ 45.519 69.961 30.907 47.659 83.084 128.667 44.068 67.348
$EWMA$ 44.704 68.490 26.626 41.728 74.557 112.995 47.751 72.910
$LW$ 42.035 65.084 24.836 40.385 75.018 113.948 43.600 66.895
$LW^{F}$ 62.001 93.989 62.649 93.590 69.286 108.383 59.458 90.055
$CCC$ 42.647 66.037 29.089 45.917 68.525 105.242 43.916 67.686
$DCC$ 40.986 63.934 26.426 42.585 69.459 106.517 42.150 65.398
$DCC^{NL}$ 42.268 65.595 27.106 43.443 71.582 109.283 43.579 67.259
$PCA$ 44.550 68.801 26.466 43.142 72.705 111.010 48.096 73.173
$RMT$ 44.334 68.575 26.291 42.936 72.803 110.768 47.761 72.940
$CAB$ 22.305 34.990 12.842 20.226 39.166 62.879 23.540 36.458
$F=90$ $NA$ 43.972 68.530 30.179 49.646 76.424 117.032 43.479 66.670
$NA^{F}$ 43.572 67.004 30.527 47.207 83.143 128.419 40.520 61.975
$EWMA$ 44.720 68.781 28.925 45.742 73.402 111.890 46.624 71.190
$LW$ 42.446 66.204 28.694 47.283 79.704 121.665 40.527 62.340
$LW^{F}$ 60.006 91.073 61.704 92.289 64.055 100.713 57.724 87.477
$CCC$ 43.390 67.342 32.735 51.757 68.495 105.076 42.999 66.475
$F=90$ $DCC$ 40.234 63.342 29.961 48.303 65.688 101.473 39.492 62.005
$DCC^{NL}$ 42.615 66.462 29.762 48.148 71.380 108.855 42.584 66.011
$PCA$ 44.464 69.171 30.520 50.091 76.867 118.267 44.085 67.265
$RMT$ 44.239 68.951 30.343 49.925 77.105 118.060 43.693 67.005
$CAB$ 13.977 21.752 10.080 16.184 20.742 32.133 14.538 22.346
$F=120$ $NA$ 48.590 75.398 36.397 59.019 84.342 128.425 46.097 70.593
$NA^{F}$ 41.909 64.465 31.713 49.127 84.615 129.958 36.093 55.352
$EWMA$ 45.330 69.861 33.556 53.195 71.953 109.497 45.222 69.142
$LW$ 47.098 73.175 34.956 56.786 87.296 132.635 43.277 66.502
$LW^{F}$ 58.216 88.420 60.114 90.286 64.569 100.736 55.133 83.621
$CCC$ 44.163 68.631 36.440 57.692 71.518 108.718 41.212 64.060
$DCC$ 40.946 64.563 33.741 54.358 68.094 104.261 37.719 59.629
$DCC^{NL}$ 44.082 68.671 34.120 54.937 74.174 112.239 41.789 64.902
$PCA$ 49.115 76.015 36.715 59.374 84.740 129.651 46.794 71.202
$RMT$ 48.867 75.827 36.519 59.220 85.041 129.486 46.352 70.985
$CAB$ 12.635 19.816 10.718 17.319 18.659 29.190 12.125 18.707
$F=180$ $NA$ 57.475 88.882 50.083 79.438 92.682 140.191 52.022 80.073
$NA^{F}$ 38.460 59.180 32.171 49.910 82.117 124.805 29.829 46.088
$EWMA$ 50.035 77.019 43.437 68.307 76.903 116.122 46.487 71.288
$LW$ 56.137 86.937 49.005 77.757 94.604 142.932 49.567 76.591
$LW^{F}$ 54.308 82.570 54.261 81.972 45.646 95.111 52.187 79.306
$CCC$ 44.949 70.232 40.531 64.242 72.901 110.150 39.678 62.498
$DCC$ 46.528 72.381 41.217 65.244 75.945 114.305 41.410 64.807
$DCC^{NL}$ 46.532 72.386 41.220 65.248 75.950 114.312 41.414 64.812
$PCA$ 58.026 89.446 50.494 79.976 92.884 141.079 52.765 80.559
$RMT$ 57.744 89.294 50.351 79.843 93.189 140.964 52.223 80.384
$CAB$ 12.356 19.146 11.116 17.602 18.159 27.468 11.472 17.726
$F=250$ $NA$ 76.850 117.44 107.237 163.541 79.505 120.16 56.341 86.706
$NA^{F}$ 34.002 52.367 33.076 51.123 64.126 97.309 25.830 40.087
$EWMA$ 53.737 82.720 53.030 82.554 72.648 109.762 48.689 74.953
$LW$ 74.903 114.697 103.984 158.986 80.825 122.029 54.291 83.797
$LW^{F}$ 49.653 75.499 47.293 71.542 45.646 70.522 52.353 79.519
$CCC$ 43.487 68.479 42.628 67.459 58.966 89.895 39.537 62.904
$DCC$ 40.299 64.494 38.892 62.869 53.375 82.622 37.405 60.269
$DCC^{NL}$ 40.223 66.604 49.514 77.717 81.610 120.697 22.135 43.633
$PCA$ 77.628 118.284 108.476 165.211 79.544 120.791 57.035 87.075
$RMT$ 77.315 118.139 108.26 165.074 79.841 120.668 56.481 86.919
$CAB$ 16.116 24.758 22.511 34.543 16.201 24.572 11.938 18.457
Table A3: Average Euclidean ($\mathcal{L}^{E}$) and Frobenius ($\mathcal{L}^{F}$) distances (×$\times 10^{5}$) using raw returns with different horizons ($F$) for different testing periods: Overall (1 January 2021 to 31 December 2023), Bull-1 (1 January 2021 to 2 January 2022), Bear (3 January 2022 to 12 June 2022), and Bull-2 (13 June 2022 to 31 December 2023).
Overall Bull-1 Bear Bull-2
$\mathcal{L}^{E}$ $\mathcal{L}^{F}$ $\mathcal{L}^{E}$ $\mathcal{L}^{F}$ $\mathcal{L}^{E}$ $\mathcal{L}^{F}$ $\mathcal{L}^{E}$ $\mathcal{L}^{F}$
$F=10$ $NA$ 70.309 107.739 54.193 82.599 107.449 165.094 69.959 107.362
$NA^{F}$ 62.057 95.297 46.340 71.028 107.068 164.102 59.155 91.020
$EWMA$ 58.908 90.338 42.125 64.612 94.704 145.253 59.384 91.053
$LW$ 63.547 97.680 47.376 72.755 104.804 160.927 62.035 95.448
$LW^{F}$ 75.854 115.361 69.181 104.007 105.831 162.355 71.457 109.048
$CCC$ 58.092 89.783 42.384 65.733 95.843 147.035 57.300 88.730
$DCC$ 62.186 94.973 48.470 73.686 96.894 147.968 60.986 93.364
$DCC^{NL}$ 65.338 99.111 51.367 77.516 100.447 152.608 64.186 97.557
$PCA$ 71.265 109.164 54.983 83.750 110.813 170.304 70.322 107.863
$RMT$ 71.064 108.951 54.818 83.587 110.625 170.014 70.093 107.640
$CAB$ 58.272 89.475 43.967 67.109 94.554 145.032 56.996 87.822
$F=40$ $NA$ 45.876 70.278 24.754 38.850 72.131 108.833 51.948 79.462
$NA^{F}$ 47.200 72.577 30.161 46.526 91.426 141.121 45.386 69.534
$EWMA$ 44.796 68.618 26.174 40.545 74.005 112.399 48.383 74.100
$LW$ 43.444 66.733 23.834 37.275 76.161 114.700 46.652 71.896
$LW^{F}$ 63.845 96.725 62.359 93.102 78.638 122.279 60.502 91.636
$CCC$ 43.321 67.023 27.469 43.276 75.123 114.858 44.354 68.515
$DCC$ 42.585 66.060 27.081 42.811 74.082 113.527 43.481 67.335
$DCC^{NL}$ 43.739 67.556 28.016 44.008 75.278 115.106 44.766 69.003
$PCA$ 46.394 71.012 24.934 39.237 72.291 109.562 52.789 80.423
$RMT$ 46.181 70.777 24.796 39.027 72.408 109.265 52.432 80.188
$CAB$ 26.619 41.417 13.427 20.765 45.197 69.510 29.777 46.649
$F=60$ $NA$ 44.060 68.125 26.115 42.593 72.109 109.679 47.545 72.605
$NA^{F}$ 45.268 69.612 30.016 46.385 84.049 130.081 43.879 67.086
$EWMA$ 44.759 68.570 26.625 41.727 74.548 112.981 47.861 73.069
$LW$ 42.043 65.095 24.836 40.385 75.020 113.951 43.615 66.916
$LW^{F}$ 62.000 93.988 62.634 93.568 69.301 108.405 59.462 90.061
$CCC$ 42.543 65.895 28.943 45.693 68.234 104.807 43.895 67.683
$DCC$ 39.404 61.909 26.764 42.978 64.589 100.135 40.279 63.072
$DCC^{NL}$ 42.240 65.544 27.015 43.295 71.326 108.855 43.658 67.380
$PCA$ 44.557 68.811 26.466 43.142 72.705 111.009 48.109 73.194
$RMT$ 44.341 68.586 26.291 42.935 72.803 110.767 47.775 72.961
$CAB$ 24.079 37.328 13.698 21.649 43.977 68.513 25.026 38.428
$F=90$ $NA$ 43.982 68.545 30.179 49.646 76.435 117.049 43.496 66.694
$NA^{F}$ 43.314 66.645 29.713 46.043 84.231 130.003 40.232 61.574
$EWMA$ 44.761 68.841 28.925 45.742 73.405 111.893 46.704 71.306
$LW$ 42.457 66.219 28.694 47.283 79.715 121.682 40.544 62.365
$LW^{F}$ 60.007 91.075 61.690 92.268 64.073 100.740 57.731 87.486
$CCC$ 43.276 67.199 32.551 51.513 68.021 104.421 43.036 66.546
$DCC$ 40.134 63.219 29.870 48.186 65.191 100.793 39.503 62.039
$DCC^{NL}$ 42.532 66.357 29.680 48.045 70.880 108.147 42.622 66.080
$PCA$ 44.474 69.186 30.520 50.092 76.878 118.284 44.100 67.288
$RMT$ 44.249 68.966 30.343 49.925 77.116 118.077 43.709 67.029
$CAB$ 14.661 23.076 10.115 16.203 22.040 34.086 15.465 24.333
$F=120$ $NA$ 48.604 75.417 36.397 59.018 84.366 128.460 46.116 70.620
$NA^{F}$ 41.664 64.124 31.048 48.181 85.754 131.620 35.717 54.821
$EWMA$ 45.361 69.908 33.555 53.194 71.959 109.504 45.282 69.232
$LW$ 47.111 73.194 34.955 56.785 87.320 132.670 43.296 66.530
$LW^{F}$ 58.219 88.425 60.101 90.267 64.599 100.781 55.138 83.629
$CCC$ 44.126 68.610 36.335 57.587 71.120 108.182 41.325 64.245
$F=120$ $DCC$ 43.988 68.569 34.032 54.862 73.636 111.512 41.819 64.965
$DCC^{NL}$ 44.025 68.619 34.051 54.887 73.690 111.582 41.864 65.025
$PCA$ 49.129 76.034 36.715 59.373 84.764 129.686 46.813 71.230
$RMT$ 48.881 75.846 36.519 59.219 85.065 129.521 46.371 71.013
$CAB$ 12.118 18.831 9.496 15.248 17.864 27.435 12.147 18.653
$F=180$ $NA$ 57.494 88.910 50.085 79.441 92.730 140.261 52.044 80.105
$NA^{F}$ 38.294 58.953 31.839 49.444 83.328 126.577 29.370 45.434
$EWMA$ 50.062 77.060 43.438 68.309 76.943 116.140 46.536 71.363
$LW$ 56.156 86.964 49.007 77.759 94.651 143.001 49.589 76.623
$LW^{F}$ 54.314 82.579 54.255 81.962 61.748 95.188 52.188 79.307
$CCC$ 44.844 70.073 40.368 64.069 72.299 109.109 39.755 62.603
$DCC$ 46.364 72.137 41.041 65.057 75.312 113.195 41.391 64.777
$DCC^{NL}$ 46.406 72.191 41.077 65.104 75.366 113.267 41.432 64.830
$PCA$ 58.045 89.474 50.496 79.979 92.932 141.148 52.787 80.592
$RMT$ 57.763 89.322 50.352 79.846 93.237 141.034 52.245 80.416
$CAB$ 12.828 19.854 11.180 17.728 20.372 30.851 11.702 18.031
$F=250$ $NA$ 76.900 117.511 107.237 163.540 79.568 120.252 56.420 86.818
$NA^{F}$ 34.000 52.376 33.222 51.351 65.353 99.107 25.373 39.432
$EWMA$ 53.792 82.799 53.034 82.561 72.673 109.796 48.785 75.090
$LW$ 74.952 114.767 103.983 158.985 80.888 122.122 54.369 83.906
$LW^{F}$ 49.692 75.555 47.295 71.544 45.710 70.616 52.408 79.599
$CCC$ 43.349 68.233 42.431 67.189 58.278 88.611 39.597 62.977
$DCC$ 40.167 64.257 38.751 62.682 52.667 81.284 37.446 60.321
$DCC^{NL}$ 40.261 66.455 49.351 77.363 80.796 119.188 22.552 44.012
$PCA$ 77.678 118.355 108.475 165.210 79.607 120.884 57.115 87.187
$RMT$ 77.365 118.210 108.260 165.073 79.904 120.760 56.560 87.032
$CAB$ 18.246 27.828 29.049 44.148 14.976 22.681 12.182 18.727
Table A4: Performance of the GMV portfolios constructed using the covariance forecasts from the 11 models under consideration plus the equal-weighted portfolio for the different market regimes. The portfolios are compared based on their annualised out-of-sample standard deviation ($\sigma_{m}$) and average out-of-sample turnover ($\text{TO}_{m}$), respectively.
Daily Weekly Monthly
$\sigma_{m}^{2}$ $TO_{m}$ $\sigma_{m}^{2}$ $TO_{m}$ $\sigma_{m}^{2}$ $TO_{m}$
Bull-1 $1/N$ 0.009883 0.005169 0.009873 0.002335 0.009907 0.001103
$NA$ 0.000543 0.150112 0.000794 0.060372 0.000834 0.026885
$NA^{F}$ 0.000731 0.001378 0.000732 0.000607 0.000736 0.000327
$EWMA$ 0.000735 0.088541 0.000748 0.043884 0.000754 0.015613
$LW$ 0.001054 0.091979 0.001572 0.044883 0.001692 0.022283
$LW^{F}$ 0.000780 0.001228 0.000781 0.000610 0.000783 0.000349
$CCC$ 0.000620 0.119486 0.000681 0.053221 0.000728 0.013605
$DCC$ 0.000529 0.207262 0.000698 0.093558 0.000781 0.024324
$DCC^{NL}$ 0.000522 0.217432 0.000784 0.097809 0.000875 0.034673
$PCA$ 0.000850 0.236835 0.000827 0.090803 0.000795 0.020906
$RMT$ 0.000856 0.150601 0.000860 0.067402 0.000826 0.025859
$CAB$ 0.000634 0.086489 0.000656 0.055931 0.000707 0.023736
Bear $1/N$ 0.029939 0.007221 0.029940 0.003443 0.029793 0.001599
$NA$ 0.002207 0.092111 0.002682 0.032018 0.002959 0.018142
$NA^{F}$ 0.002891 0.001650 0.002892 0.000935 0.002906 0.000436
$EWMA$ 0.002908 0.079065 0.002843 0.036314 0.003120 0.019997
$LW$ 0.003118 0.095723 0.003970 0.041663 0.004754 0.021610
$LW^{F}$ 0.003196 0.002035 0.003197 0.001043 0.003201 0.000570
$CCC$ 0.002497 0.129487 0.002646 0.043843 0.002590 0.012451
$DCC$ 0.002355 0.067768 0.002639 0.015372 0.002636 0.005911
$DCC^{NL}$ 0.002269 0.083567 0.002652 0.024277 0.002730 0.010308
$PCA$ 0.002800 0.114256 0.002792 0.043637 0.002864 0.011111
$RMT$ 0.002983 0.094396 0.002896 0.037508 0.003153 0.023308
$CAB$ 0.002273 0.030506 0.002464 0.022686 0.002656 0.016836
Bull-2 $1/N$ 0.019967 0.005666 0.019956 0.002382 0.019933 0.001334
$NA$ 0.002823 0.118997 0.003526 0.053457 0.003667 0.029686
$NA^{F}$ 0.003398 0.001426 0.003400 0.000577 0.003406 0.000261
$EWMA$ 0.003432 0.073812 0.003433 0.034460 0.003630 0.016887
$LW$ 0.003512 0.102531 0.004274 0.048313 0.004245 0.020765
$LW^{F}$ 0.003874 0.002113 0.003879 0.001123 0.003889 0.000788
$CCC$ 0.003233 0.129874 0.003635 0.047812 0.003671 0.015276
$DCC$ 0.002947 0.071444 0.003294 0.026871 0.003364 0.009476
$DCC^{NL}$ 0.002924 0.085242 0.003318 0.034070 0.003365 0.012661
$PCA$ 0.003793 0.194898 0.003678 0.063899 0.003597 0.024379
$RMT$ 0.003668 0.118539 0.003720 0.055498 0.003762 0.029837
$CAB$ 0.003323 0.087821 0.003547 0.052596 0.003885 0.029487

### v2

@@
 1 Introduction
-Covariance matrix estimation is central in modern finance.
+Covariance matrix forecast is central in modern finance.
 It influences portfolio construction, risk management, and asset pricing models.
 Markowitz, (1952) introduced mean-variance optimisation, highlighting how correlation among assets shapes the risk-return trade-off.
-Beyond portfolio choice, covariance estimation underpins factor-based models, hedging strategies, and even debates on market efficiency, where accurate correlation forecasts can drive risk control and potential outperformance.
+Beyond portfolio choice, covariance forecast underpins factor-based models, hedging strategies, and even debates on market efficiency, where accurate correlation forecasts can drive risk control and potential outperformance.
 Early studies found that even simple correlation structures could improve capital allocation.
@@
 (Bollerslev et al.,, 2018).
-However, GARCH-type approaches struggle with large cross-sections, falling prey to the curse of dimensionality (Robert F.
-Engle and Wolf,, 2019).
+However, GARCH-type approaches struggle with large cross-sections, falling prey to the curse of dimensionality (Engle et al.,, 2019).
 Tools like principal component analysis and random matrix theory (Ledoit and Wolf,, 2003; Laloux et al.,, 2000) have been effective for large datasets.
-Recent advances, such as dynamic conditional correlation with nonlinear shrinkage (Robert F.
-Engle and Wolf,, 2019), perform well overall, as per De Nard et al., (2024).
+Recent advances, such as dynamic conditional correlation with nonlinear shrinkage (Engle et al.,, 2019), perform well overall, as per De Nard et al., (2024).
 Still, most of these methods focus on very short-term horizons, often just one step ahead or one day ahead, and rely on high-frequency data.
-Long-horizon forecasts matter for institutional investors like pension funds and endowments, which rebalance portfolios over weeks or months.
-Moreover, high-frequency data are not always public or cost-effective, making the short-horizon focus less relevant for these institutions.
+Long-horizon forecasts matter for institutional investors, such as pension and investment funds.
+They rebalance portfolios over weeks or months.
+Therefore, a one-step forecast is insufficient for practical decision-making (De Nard et al.,, 2022).
 Furthermore, slow-moving trends, such as equity–bond correlation shifts, can be overlooked if one relies solely on high-frequency econometric models.
 Sandoval and Franca, (2012) highlights how regime shifts or significant structural changes often go undetected in classical specifications, raising the risk of large drawdowns and poor capital allocation.
-For longer-term horizons, naive, shrinkage, and factor models are more common (Ledoit and Wolf,, 2004; De Nard et al.,, 2022).
+For longer-term horizons, naive and factor models are more common (Ledoit and Wolf,, 2004; De Nard et al.,, 2022).
+Factor models can provide insightful forecasts but may not generalise well across diverse portfolios, especially when different portfolios are affected by many different factors.
+These methods are typically stable but can underperform when major shifts occur.
 Naive approaches assume that recent covariance estimates strongly predict future covariances as they follow a Markov process.
-At the same time, shrinkage techniques blend the sample covariance matrix with structured targets to reduce noise, particularly when the cross-section of assets is large (Ledoit and Wolf,, 2022).
-Factor models can provide more insight but may not generalise well across diverse portfolios, especially when markets are affected by many different factors.
-These methods are typically stable but can underperform when major shifts occur.
-In recent years, there has been a surge in machine learning methods that can handle non-linearities and high dimensionality.
+However, these tend to be noisy when the number of assets tends to be close to the number of observations ($N\approx T$).
+For this purpose, shrinkage techniques reduce noise by blending the sample covariance matrix with structured targets, particularly when the cross-section of assets is large (Ledoit and Wolf,, 2022).
+While these methods rely on advanced statistical and econometric methods, we have seen a recent surge in machine learning methods that can also handle non-linearities and high dimensionality with more accurate forecast.
 Gu et al., (2020) and Kim et al., (2023) use techniques like random forests, support vector machines, and generative adversarial networks, while Zhang et al., (2024) applies graph neural networks to capture volatility spillovers for covariance matrix forecast.
@@
 Attention mechanisms (Vaswani et al.,, 2017) offer a promising direction by weighting key observations more heavily.
-Some studies apply these techniques in finance but still focus on short-horizon tasks (Nazareth and Ramana Reddy,, 2023; Olorunnimbe and Viktor,, 2023).
+Some studies apply these techniques in finance but still focus on short-horizon tasks (Nazaret and Reddy,, 2023; Olorunnimbe and Viktor,, 2023).
 In principle, attention-based models could capture the slow-moving, cyclical factors and abrupt shifts that define medium- and long-horizon covariance forecasts, realising this potential.
@@
 Recent efforts to extend these approaches to multi-asset portfolios have been limited.
-While many studies handle large cross-sections of stocks or focus on single-asset classes such as equities (Reis et al.,, 2024), fewer examine the combined dynamics across multiple asset classes in an integrated framework.
+While many studies handle large cross-sections of stocks or focus on single-asset classes (Reis et al.,, 2024), fewer examine the combined dynamics across multiple asset classes in an integrated framework.
 However, institutional investors routinely diversify across equities, bonds, alternatives, and derivatives, underscoring the need for robust, multi-asset covariance forecasting models.
@@
 Second, there is limited empirical evidence on whether using raw returns versus excess returns leads to systematically different covariance forecasts, especially over these medium-term horizons.
-Economic theory suggests that risk premiums and the underlying drivers of asset co-movements could be more accurately captured through excess returns.
+Economic theory suggests that underlying drivers of asset co-movements could be more accurately captured through excess returns (Fama and French,, 1993), i.e.
+by removing the risk-free component from returns.
 However, most machine learning and econometric approaches mostly default to raw returns.
@@
 Specifically, we propose a novel deep learning (DL) model that captures spatio-temporal correlations while remaining robust to structural changes.
-We benchmark against classical methods (naive, shrinkage, GARCH, and dimensional reduction) and demonstrate consistent gains in forecast accuracy under diverse market environments from 2017 to 2023.
+We benchmark against classical methods and demonstrate consistent gains in forecast accuracy under diverse market environments from 2017 to 2023.
 Our main contributions are as follows:
@@
 where $\mathcal{F}_{t}$ is the sigma-algebra representing all the information available at time $t$.
-To estimate ${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}$, we compare naive and other benchmark models found in the literature alongside our proposed model. A description of each model is shown below.
-2.1 Naïve ($NA$)
+To estimate ${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}$, we compare 10 benchmark models divided into three groups (naive, de-noising and GARCH models) alongside our proposed model.
+A description of each model is shown below.
+2.1 Naïve Models
+2.1.1 Naïve ($NA$)
 A naive covariance forecasting approach is based on the lagged realised covariance.
@@
 This forecasting method is as simple as possible as it requires no optimisation.
-2.2 Naïve full sample($NA^{F}$)
+2.1.2 Naïve full sample ($NA^{F}$)
 Another naïve technique is to use the full sample instead of the rolling covariance matrix as our estimator.
@@
 where $\mathbf{\Sigma}$ is the covariance matrix of our sample up until $t$.
-2.3 Exponential Weighted Moving Average ($EWMA$)
+2.1.3 Exponential Weighted Moving Average ($EWMA$)
 The EWMA employs exponentially decaying weights for the covariance matrix. The covariance matrix in the EWMA model is recursively computed as follows:
@@
 This factor has been approximated to be around 0.94 (J.P.Morgan,, 1996).
-2.4 Ledoit and Wolf ($LW$)
-The LW shrinkage (Ledoit and Wolf,, 2004) is a robust statistical model that combines the sample covariance matrix with a structured target matrix, improving estimation accuracy by reducing the impact of estimation error due to high dimensionality and limited sample size.
+2.2 De-noising Models
+2.2.1 Principal Component Analysis ($PCA$)
+As noted by Ledoit and Wolf, (2003), empirical results show the usefulness of PCA in covariance matrix forecast. The PCA is computed as follows:
+${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}=\mathbf{U\Lambda}\mathbf{U}^{\top}$ (10)
+where $\mathbf{U}$ is the matrix of eigenvectors and $\mathbf{\Lambda}=\text{diag}(\lambda_{1},\lambda_{2},...,\lambda_{N})$ is the diagonal matrix of eigenvalues.
+To ensure we capture most of the information, we retain only the first $k$ principal components that account for at least 95% of the total variance. This is expressed as:
+$\sum_{i=1}^{k}\lambda_{i}\geq 0.95\sum_{i=1}^{N}\lambda_{i}$ (11)
+The final forecasted covariance matrix is then approximated as:
+${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}\approx\mathbf{U}_{k}\mathbf{\Lambda}_{k}\mathbf{U}_{k}^{\top}$ (12)
+2.2.2 Random Matrix Theory ($RMT$)
+$RMT$ (Laloux et al.,, 2000) is built upon the $PCA$ where the upper ($\lambda_{+}$) and lower ($\lambda_{-}$) bound of the eigenvalues are calculated using the Marchenko-Pastur distribution.
+These bounds are calculated as:
+$\lambda_{+}=\left(1+\frac{1}{\sqrt{q}}\right)^{2},\quad\lambda_{-}=\left(1-\frac{1}{\sqrt{q}}\right)^{2}$ (13)
+where $q=T/N$ is the ratio of the window size to the number of assets. Eigenvalues within these bounds are dominated by noise, while eigenvalues outside the bounds represent meaningful information.
+$\lambda_{i}^{\text{filtered}}=\begin{cases}\lambda_{i}&\text{if }\lambda_{i}>\lambda_{+}\text{ or }\lambda_{i}<\lambda_{-}\\ \frac{\lambda_{+}+\lambda_{-}}{2}&\text{otherwise}\end{cases}$ (14)
+This filtering process removes noise and retains only the significant components of the covariance matrix.
+The cleaned covariance matrix is reconstructed by multiplying the filtered eigenvalues ($\mathbf{\Lambda}^{\text{filtered}}$) with their corresponding eigenvectors:
+${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}=\mathbf{U}\mathbf{\Lambda}^{\text{filtered}}\mathbf{U}^{\top}$ (15)
+2.2.3 Ledoit and Wolf ($LW$)
+The LW shrinkage (Ledoit and Wolf,, 2004) is a robust statistical model that combines the sample covariance matrix with a structured target matrix, improving forecast accuracy by reducing the impact of forecast error due to high dimensionality and limited sample size.
 The LW shrinkage estimator is given by:
@@
 where $\sigma_{i,j}$ are the elements of the sample covariance matrix and $t_{i,j}$ are the elements of the target matrix $\mathbf{T}$.
-2.5 Ledoit and Wolf full sample ($LW^{F}$)
+2.2.4 Ledoit and Wolf full sample ($LW^{F}$)
 Similarly to the $NA^{F}$ model, we employed a full sample of Ledoit and Wolf, (2004) as a shrinkage model.
 ${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}=\rho\mathbf{T}+(1-\rho)\mathbf{\Sigma}$ (18)
-2.6 Constant Conditional Correlations ($CCC$)
+2.3 GARCH Models
+2.3.1 Constant Conditional Correlations ($CCC$)
 The $CCC$ model, introduced by Bollerslev, (1990), assumes that asset correlations remain constant over time while variances evolve dynamically according to a GARCH process based on a one-step forecast.
-However, since investment funds and risk managers typically rebalance portfolios over extended horizons, such as monthly or quarterly, a one-step forecast is insufficient for practical decision-making (De Nard et al.,, 2022).
 To extend the model to a multi-step forecasting framework, we follow the methodology of Baillie and Bollerslev, (1992), defining the $F$-step ahead conditional covariance forecast as:
@@
 where $e_{i,t}$ are the residuals of the return series, which are used to estimate the conditional correlation matrix.
-2.7 Dynamic Conditional Correlations ($DCC$)
+2.3.2 Dynamic Conditional Correlations ($DCC$)
 The $DCC$ model of Engle, (2002) extends the $CCC$ model by allowing correlations to vary over time while retaining the GARCH framework for variances.
@@
 (21), and $\overline{\mathbf{Q}}$ is estimated from historical standardised residuals.
-2.8 Nonlinear Shrinkage Dynamic Conditional Correlations ($DCC^{NL}$)
-One recent extension is the $DCC^{NL}$ model from De Nard et al., (2022).
+2.3.3 Nonlinear Shrinkage Dynamic Conditional Correlations ($DCC^{NL}$)
+One work that combines the multivariate GARCH with a nonlinear shrinkage is the model of Engle et al., (2019).
 The covariance forecast in the $DCC^{NL}$ model follows the same structure as in Eq.
 (22), but now with a shrinkage-improved correlation matrix ($\overline{\mathbf{Q}}$).
-They apply the nonlinear shrinkage method of Ledoit and Wolf, (2020), which optimally shrinks the eigenvalues of the sample covariance matrix towards a better-conditioned target matrix.
-This procedure ensures that the estimated correlation matrix remains positive, definite, and well-conditioned, even in high-dimensional settings.
+The original model uses the numerical method of Ledoit and Wolf, (2015).
+However, we apply the nonlinear shrinkage method of Ledoit and Wolf, (2020) to speed up the computations, which optimally shrinks the eigenvalues of the sample covariance matrix towards a better-conditioned target matrix.
 Given a sample correlation matrix $\mathbf{\Upsilon}$, the optimal nonlinear shrinkage estimator adjusts its eigenvalues as follows. Let $\mathbf{\Upsilon}$ have the eigendecomposition:
 $\mathbf{\Upsilon}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top},$ (26)
-where $\mathbf{U}$ is the matrix of eigenvectors and $\mathbf{\Lambda}=\text{diag}(\lambda_{1},\lambda_{2},...,\lambda_{N})$ is the diagonal matrix of eigenvalues.
 The nonlinear shrinkage transformation replaces $\mathbf{\Lambda}$ with a shrinkage-transformed version $\mathbf{\Lambda}^{*}$:
@@
 The shrunk correlation matrix ($\mathbf{\Upsilon}^{*}$) is then used as the input for the $DCC$ process in Eq. (24).
-2.9 Principal Component Analysis ($PCA$)
-As noted by Ledoit and Wolf, (2003), empirical results show the usefulness of PCA in covariance matrix estimation. The PCA is computed as follows:
-${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}=\mathbf{U\Lambda}\mathbf{U}^{\top}$ (23)
-To ensure we capture most of the information, we retain only the first $k$ principal components that account for at least 95% of the total variance. This is mathematically expressed as:
-$\sum_{i=1}^{k}\lambda_{i}\geq 0.95\sum_{i=1}^{N}\lambda_{i}$ (24)
-The final forecasted covariance matrix is then approximated as:
-${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}\approx\mathbf{U}_{k}\mathbf{\Lambda}_{k}\mathbf{U}_{k}^{\top}$ (25)
-2.10 Random Matrix Theory ($RMT$)
-$RMT$ (Laloux et al.,, 2000) is built upon the $PCA$ where the upper ($\varsigma_{+}$) and lower ($\varsigma_{-}$) bound of the eigenvalues are calculated using the Marchenko-Pastur distribution.
-These bounds are calculated as:
-$\lambda_{+}=\left(1+\frac{1}{\sqrt{q}}\right)^{2},\quad\lambda_{-}=\left(1-\frac{1}{\sqrt{q}}\right)^{2}$ (26)
-where $q=T/N$ is the ratio of the window size to the number of assets. Eigenvalues within these bounds are dominated by noise, while eigenvalues outside the bounds represent meaningful information.
-$\lambda_{i}^{\text{filtered}}=\begin{cases}\lambda_{i}&\text{if }\lambda_{i}>\lambda_{+}\text{ or }\lambda_{i}<\lambda_{-}\\ \frac{\lambda_{+}+\lambda_{-}}{2}&\text{otherwise}\end{cases}$ (27)
-This filtering process removes noise and retains only the significant components of the covariance matrix.
-The cleaned covariance matrix is reconstructed by multiplying the filtered eigenvalues with their corresponding eigenvectors:
-${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}=\mathbf{U}\mathbf{\Lambda}^{\text{filtered}}\mathbf{U}^{\top}$ (28)
 2.4 CNN with A-BiLSTM ($CAB$)
@@
 However, in our model, we use a similar approach to Liu and Guo, (2019) methodology, meaning all time steps from the last layer ($\mathbf{H}^{u}$) are used as an input to our following steps.
+This procedure allows the attention mechanism to dynamically weigh the importance of each time step, enabling the model to focus on the most relevant temporal patterns when making predictions.
+By leveraging all time steps from the last layer ($\mathbf{H}^{u}$), rather than just the final output, the model can more effectively capture long-term dependencies and contextual relationships.
 To mitigate overfitting and enhance the model’s generalisation capabilities, a Dropout Layer is applied to the BiLSTM output.
@@
 2.4.6 Shrinkage
-While DL-based forecasts can learn intricate patterns, historical covariance may still convey valuable stability and continuity, particularly when the model is tested on data different from training.
+While DL-based forecasts can learn intricate patterns, which provide a more stable forecast, the noisy historical covariance may still convey valuable information about recent market conditions.
 Therefore, our model’s final step is incorporating a linear shrinkage from historical values in our predictions.
@@
 Finally, we apply a shrinkage factor of 0.8 ($\phi$) when combining the DL output with the most recent historical covariance.
-We conduct a rolling estimation procedure to simulate live conditions.
+We conduct a rolling forecast procedure to simulate live conditions.
 Each new trading day $t$ in the test set triggers an online update of our DL model, incorporating all data up through day $t$.
-This design ensures that the forecasts mimic a real-world trading environment, where only up-to-date historical data is available at each decision point.
+This design ensures the forecasts mimic a real-world trading environment, where only up-to-date historical data is available at each decision point.
 Table 2 reports the average Euclidean ($\mathcal{L}^{E}$) and Frobenius ($\mathcal{L}^{F}$) distances between forecasted and realised covariance matrices for all models over the entire testing period (2021–2023) [^4] Bold and underlined numbers represent the lowest and the second lowest values, respectively..
 GARCH models ($CCC$ and $DCC$) and $EWMA$ perform relatively well as they adapt faster to structural changes in the covariance matrices.
+Please note that although $DCC^{NL}$ is not more accurate than the traditional $DCC$, it provides more stable forecasts based on the eigenvalues.
 However, our model ($CAB$) yields the lowest errors in both metrics, outperforming classical benchmarks.
@@
 $EWMA$ 49.111 75.349
+$PCA$ 54.348 83.154
+$RMT$ 54.145 82.915
 $LW$ 49.772 76.514
@@
 $DCC^{NL}$ 51.621 79.079
-$PCA$ 54.348 83.154
-$RMT$ 54.145 82.915
 $CAB$ 38.312 58.923
@@
 $EWMA$ 31.754 48.973 82.799 126.614 50.574 77.549
+$PCA$ 37.424 57.281 90.720 139.116 54.747 83.660
+$RMT$ 37.246 57.083 90.638 138.821 54.493 83.411
 $LW$ 34.126 52.599 90.441 138.173 48.089 74.088
@@
 $DCC^{NL}$ 36.779 56.449 87.650 134.167 50.767 77.734
-$PCA$ 37.424 57.281 90.720 139.116 54.747 83.660
-$RMT$ 37.246 57.083 90.638 138.821 54.493 83.411
 $CAB$ 24.276 37.643 66.381 102.252 39.253 60.126
@@
 4 Robustness analysis
+4.1 Larger dataset
+In this section, we expand our asset universe to 260 ETFs to evaluate whether our model continues to outperform established benchmarks in a high-dimensional setting.
+Appendix LABEL:tab:listofetfslarge provides a comprehensive list of the ETFs in a multi-asset portfolio.
+Table 4 presents the corresponding performance metrics.
+Table 4: Average Euclidean ($\mathcal{L}^{E}$) and Frobenius ($\mathcal{L}^{F}$) distances (×$\times 10^{3}$) using 260 ETFs for $F=20$ and excess returns for different testing periods: Overall (1 January 2021 to 31 December 2023), Bull-1 (1 January 2021 to 2 January 2022), Bear (3 January 2022 to 12 June 2022), and Bull-2 (13 June 2022 to 31 December 2023).
+Overall Bull-1 Bear Bull-2
+$\mathcal{L}^{E}$ $\mathcal{L}^{F}$ $\mathcal{L}^{E}$ $\mathcal{L}^{F}$ $\mathcal{L}^{E}$ $\mathcal{L}^{F}$ $\mathcal{L}^{E}$ $\mathcal{L}^{F}$
+$NA$ 11.546 16.484 8.677 12.442 20.458 29.257 10.814 15.390
+$NA^{F}$ 10.792 15.399 7.661 11.014 20.742 29.532 9.927 14.130
+$EWMA$ 10.544 15.054 7.518 10.802 18.410 26.292 10.219 14.543
+$LW$ 10.719 15.304 8.095 11.619 19.933 28.458 9.740 13.867
+$LW^{F}$ 13.460 19.170 12.098 17.270 19.374 27.609 12.622 17.947
+$CCC$ 10.260 14.647 7.368 10.587 18.723 26.687 9.674 13.777
+$DCC$ 10.280 14.671 7.796 11.182 18.076 25.777 9.622 13.703
+$DCC^{NL}$ 11.047 15.747 8.597 12.301 18.687 26.634 10.414 14.814
+$PCA$ 11.658 16.643 8.743 12.535 20.625 29.500 10.939 15.567
+$RMT$ 11.652 16.636 8.738 12.529 20.618 29.490 10.934 15.561
+$CAB$ 8.933 12.789 5.447 7.893 15.722 22.456 9.220 13.153
+Our proposed approach ($CAB$) consistently outperforms the benchmark models, reflected by its lower average Euclidean and Frobenius distances across most periods.
+Notably, during the Overall, Bull-1, and Bull-2 periods, the post-hoc Nemenyi test confirms that the outperformance of CAB is statistically significant.
+In contrast, during the Bear period, CAB’s performance is statistically similar to that of the $DCC$ and $EWMA$ models.
 4.2 Impact of Using Raw Returns vs. Excess Returns
@@
 When comparing Table 5 (raw returns) to Tables 2 and 3 (excess returns), the numerical magnitudes change slightly, but the rank order remains the same: $CAB$ still produces the most accurate forecasts overall.
-Our findings indicate that the advantage of our framework is not contingent on the choice between excess or raw returns, reaffirming its robustness to the presence or absence of the risk-free rate.
-Moreover, this consistency suggests that subtracting the risk-free rate over the medium-term horizon does not significantly alter the core structure of asset co-movements.
-Hence, even though economic theory posits that excess returns better capture risk premiums, our empirical results show that the forecasting edge of $CAB$ holds across both raw and excess return specifications.
-In other words, our method retains its superior performance even when modelling raw returns, suggesting that its effectiveness does not hinge on subtracting the risk-free rate.
 Table 5: Average Euclidean ($\mathcal{L}^{E}$) and Frobenius ($\mathcal{L}^{F}$) distances (×$\times 10^{5}$) using raw returns for different testing periods: Overall (1 January 2021 to 31 December 2023), Bull-1 (1 January 2021 to 2 January 2022), Bear (3 January 2022 to 12 June 2022), and Bull-2 (13 June 2022 to 31 December 2023).
@@
 $EWMA$ 49.174 75.441 31.753 48.972 82.798 126.613 50.697 77.730
+$PCA$ 54.348 83.153 37.423 57.280 90.724 139.123 54.745 83.657
+$RMT$ 54.144 82.914 37.245 57.081 90.643 138.829 54.491 83.408
 $LW$ 49.771 76.512 34.125 52.598 90.442 138.174 48.087 74.085
@@
 $DCC^{NL}$ 51.604 79.051 36.691 56.305 87.856 134.535 50.731 77.664
-$PCA$ 54.348 83.153 37.423 57.280 90.724 139.123 54.745 83.657
-$RMT$ 54.144 82.914 37.245 57.081 90.643 138.829 54.491 83.408
 $CAB$ 36.663 56.518 22.807 35.673 62.036 96.406 38.273 58.441
+Our findings indicate that the advantage of our framework is not contingent on the choice between excess or raw returns, reaffirming its robustness to the presence or absence of the risk-free rate.
+Moreover, this consistency suggests that subtracting the risk-free rate over the medium-term horizon does not significantly alter the core structure of asset co-movements.
+Hence, even though economic theory posits that excess returns better capture risk premiums, our empirical results show that the forecasting edge of $CAB$ holds across both raw and excess return specifications.
+In other words, our method retains its superior performance even when modelling raw returns, suggesting that its effectiveness does not hinge on subtracting the risk-free rate.
 4.3 Varying Forecast Horizons
@@
 Excluding expected returns from the optimisation allows us to isolate the impact of covariance forecasts (DeMiguel et al.,, 2009; Kourtis et al.,, 2012).
-Using a rolling estimation window, we generate 11 portfolio strategies (one for each forecast model) plus an equally weighted (1/$N$) benchmark.
+Using a rolling forecast window, we generate 11 portfolio strategies (one for each forecast model) plus an equally weighted (1/$N$) benchmark.
 After estimating $\widehat{\mathbf{\Sigma}}_{t}$, we compute the GMV weights and hold them until the next rebalancing date.
@@
 $EWMA$ 0.002475 0.079388 0.002482 0.037928 0.002554 0.016088
+$PCA$ 0.002680 0.197322 0.002722 0.072903 0.002657 0.023108
+$RMT$ 0.002647 0.125512 0.002689 0.055954 0.002733 0.024767
 $LW$ 0.002655 0.098020 0.003373 0.047089 0.003598 0.022957
@@
 $DCC^{NL}$ 0.002048 0.129453 0.002448 0.056173 0.002618 0.023470
-$PCA$ 0.002680 0.197322 0.002722 0.072903 0.002657 0.023108
-$RMT$ 0.002647 0.125512 0.002689 0.055954 0.002733 0.024767
 $CAB$ 0.002293 0.078799 0.002428 0.048383 0.002579 0.023807
 When compared to the $1/N$ benchmark, please note that all models present a statistically significant lower annual variance, based on the one-side F-Test of equalities of variance.
+Table 7: Performance of the GMV portfolios constructed using the covariance forecasts from the 11 models under consideration plus the equal-weighted portfolio for the different market regimes.
+The portfolios are compared based on their annualised out-of-sample standard deviation ($\sigma_{m}$) and average out-of-sample turnover ($\text{TO}_{m}$), respectively.
+Daily Weekly Monthly
+$\sigma_{m}^{2}$ $TO_{m}$ $\sigma_{m}^{2}$ $TO_{m}$ $\sigma_{m}^{2}$ $TO_{m}$
+Bull-1 $1/N$ 0.009883 0.005169 0.009873 0.002335 0.009907 0.001103
+$NA$ 0.000543 0.150112 0.000794 0.060372 0.000834 0.026885
+$NA^{F}$ 0.000731 0.001378 0.000732 0.000607 0.000736 0.000327
+$EWMA$ 0.000735 0.088541 0.000748 0.043884 0.000754 0.015613
+$PCA$ 0.000850 0.236835 0.000827 0.090803 0.000795 0.020906
+$RMT$ 0.000856 0.150601 0.000860 0.067402 0.000826 0.025859
+$LW$ 0.001054 0.091979 0.001572 0.044883 0.001692 0.022283
+$LW^{F}$ 0.000780 0.001228 0.000781 0.000610 0.000783 0.000349
+$CCC$ 0.000620 0.119486 0.000681 0.053221 0.000728 0.013605
+$DCC$ 0.000529 0.207262 0.000698 0.093558 0.000781 0.024324
+$DCC^{NL}$ 0.000522 0.217432 0.000784 0.097809 0.000875 0.034673
+$CAB$ 0.000634 0.086489 0.000656 0.055931 0.000707 0.023736
+Bear $1/N$ 0.029939 0.007221 0.029940 0.003443 0.029793 0.001599
+$NA$ 0.002207 0.092111 0.002682 0.032018 0.002959 0.018142
+$NA^{F}$ 0.002891 0.001650 0.002892 0.000935 0.002906 0.000436
+$EWMA$ 0.002908 0.079065 0.002843 0.036314 0.003120 0.019997
+$PCA$ 0.002800 0.114256 0.002792 0.043637 0.002864 0.011111
+$RMT$ 0.002983 0.094396 0.002896 0.037508 0.003153 0.023308
+$LW$ 0.003118 0.095723 0.003970 0.041663 0.004754 0.021610
+$LW^{F}$ 0.003196 0.002035 0.003197 0.001043 0.003201 0.000570
+$CCC$ 0.002497 0.129487 0.002646 0.043843 0.002590 0.012451
+$DCC$ 0.002355 0.067768 0.002639 0.015372 0.002636 0.005911
+$DCC^{NL}$ 0.002269 0.083567 0.002652 0.024277 0.002730 0.010308
+$CAB$ 0.002273 0.030506 0.002464 0.022686 0.002656 0.016836
+Bull-2 $1/N$ 0.019967 0.005666 0.019956 0.002382 0.019933 0.001334
+$NA$ 0.002823 0.118997 0.003526 0.053457 0.003667 0.029686
+$NA^{F}$ 0.003398 0.001426 0.003400 0.000577 0.003406 0.000261
+$EWMA$ 0.003432 0.073812 0.003433 0.034460 0.003630 0.016887
+$PCA$ 0.003793 0.194898 0.003678 0.063899 0.003597 0.024379
+$RMT$ 0.003668 0.118539 0.003720 0.055498 0.003762 0.029837
+$LW$ 0.003512 0.102531 0.004274 0.048313 0.004245 0.020765
+$LW^{F}$ 0.003874 0.002113 0.003879 0.001123 0.003889 0.000788
+$CCC$ 0.003233 0.129874 0.003635 0.047812 0.003671 0.015276
+$DCC$ 0.002947 0.071444 0.003294 0.026871 0.003364 0.009476
+$DCC^{NL}$ 0.002924 0.085242 0.003318 0.034070 0.003365 0.012661
+$CAB$ 0.003323 0.087821 0.003547 0.052596 0.003885 0.029487
 Examining the same metrics across distinct market regimes (Table LABEL:tab:portfoliovariancesmr) reveals that $CAB$ remains a robust contender under varying conditions: in Bull-1, it consistently competes for the lowest variance, even attaining the best performance in weekly and monthly rebalancing; in the Bear regime, $CAB$ ranks near the top in daily and weekly horizons, balancing lower turnover against marginally higher volatility than a few dynamic methods; and while Bull-2 sees other estimators occasionally surpass $CAB$, it still demonstrates stable risk control with moderate trading activity, indicating that its effectiveness persists across both tranquil and turbulent market environments.
@@
 This stability benefits investors who must limit transaction costs or adhere to regulatory constraints on turnover.
-These results underscore that advanced covariance estimation can yield economically meaningful improvements in risk control, particularly over the medium-term horizons favoured by many institutional investors.
+These results underscore that advanced covariance forecast can yield economically meaningful improvements in risk control, particularly over the medium-term horizons favoured by many institutional investors.
 The $CAB$ model stands out by combining strong variance reduction with stable rebalancing behaviour, balancing responsiveness to market changes and turnover efficiency.
@@
 Notably, its performance advantage remains robust whether one subtracts a risk-free rate from returns, suggesting broad applicability in diverse portfolio management practices.
-In portfolio experiments, the proposed method enables global minimum-variance (GMV) strategies to achieve lower out-of-sample volatility with moderate turnover, underlining the tangible economic value of improved covariance estimation.
+In portfolio experiments, the proposed method enables GMV strategies to achieve lower out-of-sample volatility with moderate turnover, underlining the tangible economic value of improved covariance forecast.
 By bridging cutting-edge DL techniques with established financial principles, our work highlights the promise of sophisticated spatiotemporal modelling for risk management and allocation decisions, especially at horizons where structural shifts and evolving correlations pose unique forecasting challenges.
-Still, several research avenues warrant further exploration.
+Still, two research avenues warrant further exploration.
 First, extending the network architecture to accommodate skewed or heavy-tailed distributions could better capture tail risk in times of market stress.
 Second, explicitly incorporating transaction costs or liquidity constraints within the optimisation process might yield more realistic and implementable trading strategies.
-Finally, evaluating the model in larger cross-sectional settings or applied to alternative asset classes (e.g., commodities, crypto assets) would illuminate its scalability and robustness.
 We hope these findings encourage the finance community to explore and adapt advanced neural architectures for medium-term and long-horizon applications, strengthening the link between machine learning innovation and effective real-world asset allocation.
@@
 Shrinkage factor ($\phi$) 0.00, 0.20, 0.40, 0.60, 0.80, 1.00
-A.3 Extended Data Tables
+A.3 Extended ETF List
+Table A2: List of ETFs used in the larger dataset.
+Name Ticker Incept. Date Asset Class
+iShares Core International Aggregate Bond ETF IAGG 10/11/2015 Fixed Income
+iShares MSCI Saudi Arabia ETF KSA 16/09/2015 Equity
+iShares Interest Rate Hedged Long-Term Corporate Bond ETF IGBH 22/07/2015 Fixed Income
+iShares Currency Hedged MSCI ACWI ex U.S. ETF HAWX 29/06/2015 Equity
+iShares Currency Hedged MSCI EAFE Small-Cap ETF HSCZ 29/06/2015 Equity
+iShares MSCI Intl Value Factor ETF IVLU 16/06/2015 Equity
+iShares Convertible Bond ETF ICVT 02/06/2015 Fixed Income
+iShares U.S. Equity Factor ETF LRGF 28/04/2015 Equity
+iShares U.S. Small-Cap Equity Factor ETF SMLF 28/04/2015 Equity
+iShares International Equity Factor ETF INTF 28/04/2015 Equity
+iShares International Small-Cap Equity Factor ETF ISCF 28/04/2015 Equity
+iShares Global Equity Factor ETF GLOF 28/04/2015 Equity
+iShares Exponential Technologies ETF XT 19/03/2015 Equity
+iShares iBonds Dec 2025 Term Corporate ETF IBDQ 11/03/2015 Fixed Income
+iShares Short Maturity Municipal Bond Active ETF MEAR 03/03/2015 Fixed Income
+iShares U.S. Fixed Income Balanced Risk Systematic ETF FIBR 24/02/2015 Fixed Income
+iShares MSCI Intl Quality Factor ETF IQLT 13/01/2015 Equity
+iShares MSCI Intl Momentum Factor ETF IMTM 13/01/2015 Equity
+iShares MSCI ACWI Low Carbon Target ETF CRBN 08/12/2014 Equity
+iShares GSCI Commodity Dynamic Roll Strategy ETF COMT 15/10/2014 Commodity
+iShares Currency Hedged MSCI Emerging Markets ETF HEEM 23/09/2014 Equity
+iShares Currency Hedged MSCI Eurozone ETF HEZU 09/07/2014 Equity
+iShares Global REIT ETF REET 08/07/2014 Real Estate
+iShares Core Total USD Bond Market ETF IUSB 10/06/2014 Fixed Income
+iShares Core Dividend Growth ETF DGRO 10/06/2014 Equity
+iShares Core MSCI Europe ETF IEUR 10/06/2014 Equity
+iShares Core MSCI Pacific ETF IPAC 10/06/2014 Equity
+iShares Interest Rate Hedged Corporate Bond ETF LQDH 27/05/2014 Fixed Income
+iShares Interest Rate Hedged High Yield Bond ETF HYGH 27/05/2014 Fixed Income
+iShares MSCI UAE ETF UAE 29/04/2014 Equity
+iShares MSCI Qatar ETF QAT 29/04/2014 Equity
+iShares Yield Optimized Bond ETF BYLD 22/04/2014 Fixed Income
+iShares Treasury Floating Rate Bond ETF TFLO 03/02/2014 Fixed Income
+iShares Currency Hedged MSCI EAFE ETF HEFA 31/01/2014 Equity
+iShares Currency Hedged MSCI Japan ETF HEWJ 31/01/2014 Equity
+iShares Ultra Short-Term Bond Active ETF ICSH 11/12/2013 Fixed Income
+iShares 0-5 Year High Yield Corporate Bond ETF SHYG 15/10/2013 Fixed Income
+iShares 0-5 Year Investment Grade Corporate Bond ETF SLQD 15/10/2013 Fixed Income
+iShares Short Duration Bond Active ETF NEAR 25/09/2013 Fixed Income
+iShares MSCI USA Quality Factor ETF QUAL 16/07/2013 Equity
+iShares MSCI USA Momentum Factor ETF MTUM 16/04/2013 Equity
+iShares MSCI USA Value Factor ETF VLUE 16/04/2013 Equity
+iShares MSCI USA Size Factor ETF SIZE 16/04/2013 Equity
+iShares Core MSCI EAFE ETF IEFA 18/10/2012 Equity
+iShares Core MSCI Emerging Markets ETF IEMG 18/10/2012 Equity
+iShares Core MSCI Total International Stock ETF IXUS 18/10/2012 Equity
+iShares Core 1-5 Year USD Bond ETF ISTB 18/10/2012 Fixed Income
+iShares J.P. Morgan EM Corporate Bond ETF CEMB 17/04/2012 Fixed Income
+iShares J.P. Morgan EM High Yield Bond ETF EMHY 03/04/2012 Fixed Income
+iShares US & Intl High Yield Corp Bond ETF GHYG 03/04/2012 Fixed Income
+iShares Morningstar Multi-Asset Income ETF IYLD 03/04/2012 Multi Asset
+iShares International High Yield Bond ETF HYXU 03/04/2012 Fixed Income
+iShares Emerging Markets Dividend ETF DVYE 23/02/2012 Equity
+iShares Asia/Pacific Dividend ETF DVYA 23/02/2012 Equity
+iShares U.S. Treasury Bond ETF GOVT 14/02/2012 Fixed Income
+iShares Aaa - A Rated Corporate Bond ETF QLTA 14/02/2012 Fixed Income
+iShares CMBS ETF CMBS 14/02/2012 Others
+iShares GNMA Bond ETF GNMA 14/02/2012 Fixed Income
+iShares MSCI India Small-Cap ETF SMIN 08/02/2012 Equity
+iShares MSCI Emerging Markets Asia ETF EEMA 08/02/2012 Equity
+iShares MSCI India ETF INDA 02/02/2012 Equity
+iShares MSCI Global Gold Miners ETF RING 31/01/2012 Equity
+iShares MSCI Global Metals & Mining Producers ETF PICK 31/01/2012 Equity
+iShares MSCI Global Silver and Metals Miners ETF SLVP 31/01/2012 Equity
+iShares MSCI Agriculture Producers ETF VEGI 31/01/2012 Equity
+iShares MSCI Global Energy Producers ETF FILL 31/01/2012 Equity
+iShares MSCI Denmark ETF EDEN 25/01/2012 Equity
+iShares MSCI United Kingdom Small-Cap ETF EWUS 25/01/2012 Equity
+iShares MSCI Finland ETF EFNL 25/01/2012 Equity
+iShares MSCI Norway ETF ENOR 23/01/2012 Equity
+iShares MSCI World ETF URTH 10/01/2012 Equity
+iShares MSCI USA Min Vol Factor ETF USMV 18/10/2011 Equity
+iShares MSCI EAFE Min Vol Factor ETF EFAV 18/10/2011 Equity
+iShares MSCI Global Min Vol Factor ETF ACWV 18/10/2011 Equity
+iShares MSCI Emerging Markets Min Vol Factor ETF EEMV 18/10/2011 Equity
+iShares J.P. Morgan EM Local Currency Bond ETF LEMB 18/10/2011 Fixed Income
+iShares MSCI Emerging Markets Small-Cap ETF EEMS 16/08/2011 Equity
+iShares Floating Rate Bond ETF FLOT 14/06/2011 Fixed Income
+iShares Core High Dividend ETF HDV 29/03/2011 Equity
+iShares MSCI China ETF MCHI 29/03/2011 Equity
+iShares 0-5 Year TIPS Bond ETF STIP 01/12/2010 Fixed Income
+iShares MSCI Brazil Small-Cap ETF EWZS 28/09/2010 Equity
+iShares MSCI Philippines ETF EPHE 28/09/2010 Equity
+iShares MSCI China Small-Cap ETF ECNS 28/09/2010 Equity
+iShares MSCI New Zealand ETF ENZL 01/09/2010 Equity
+iShares MSCI Poland ETF EPOL 25/05/2010 Equity
+iShares MSCI USA Equal Weighted ETF EUSA 05/05/2010 Equity
+iShares MSCI Indonesia ETF EIDO 05/05/2010 Equity
+iShares MSCI Ireland ETF EIRL 05/05/2010 Equity
+iShares MSCI Europe Financials ETF EUFN 20/01/2010 Equity
+iShares 10+ Year Investment Grade Corporate Bond ETF IGLB 08/12/2009 Fixed Income
+iShares Core 10+ Year USD Bond ETF ILTB 08/12/2009 Fixed Income
+iShares India 50 ETF INDY 18/11/2009 Equity
+iShares Russell Top 200 Growth ETF IWY 22/09/2009 Equity
+iShares Russell Top 200 Value ETF IWX 22/09/2009 Equity
+iShares Russell Top 200 ETF IWL 22/09/2009 Equity
+iShares MSCI Peru and Global Exposure ETF EPU 19/06/2009 Equity
+iShares Emerging Markets Infrastructure ETF EMIF 16/06/2009 Equity
+iShares International Treasury Bond ETF IGOV 21/01/2009 Fixed Income
+iShares 1-3 Year International Treasury Bond ETF ISHG 21/01/2009 Fixed Income
+iShares Short-Term National Muni Bond ETF SUB 05/11/2008 Fixed Income
+iShares Agency Bond ETF AGZ 05/11/2008 Fixed Income
+iShares Core Growth Allocation ETF AOR 04/11/2008 Multi Asset
+iShares Core Aggressive Allocation ETF AOA 04/11/2008 Multi Asset
+iShares Core Moderate Allocation ETF AOM 04/11/2008 Multi Asset
+iShares Core Conservative Allocation ETF AOK 04/11/2008 Multi Asset
+iShares MSCI All Country Asia ex Japan ETF AAXJ 13/08/2008 Equity
+iShares Global Clean Energy ETF ICLN 24/06/2008 Equity
+iShares Global Timber & Forestry ETF WOOD 24/06/2008 Equity
+iShares MSCI ACWI ETF ACWI 26/03/2008 Equity
+iShares MSCI ACWI ex U.S. ETF ACWX 26/03/2008 Equity
+iShares MSCI Israel ETF EIS 26/03/2008 Equity
+iShares MSCI Thailand ETF THD 26/03/2008 Equity
+iShares MSCI Turkey ETF TUR 26/03/2008 Equity
+iShares MSCI Japan Small-Cap ETF SCJ 20/12/2007 Equity
+iShares J.P. Morgan USD Emerging Markets Bond ETF EMB 17/12/2007 Fixed Income
+iShares MSCI EAFE Small-Cap ETF SCZ 10/12/2007 Equity
+iShares Global Infrastructure ETF IGF 10/12/2007 Equity
+iShares MSCI Kokusai ETF TOK 10/12/2007 Equity
+iShares Asia 50 ETF AIA 13/11/2007 Equity
+iShares MSCI Chile ETF ECH 12/11/2007 Equity
+iShares International Developed Real Estate ETF IFGL 12/11/2007 Real Estate
+iShares MSCI Europe Small-Cap ETF IEUS 12/11/2007 Equity
+iShares MSCI BIC ETF BKF 12/11/2007 Equity
+iShares California Muni Bond ETF CMF 04/10/2007 Fixed Income
+iShares New York Muni Bond ETF NYF 04/10/2007 Fixed Income
+iShares National Muni Bond ETF MUB 07/09/2007 Fixed Income
+iShares International Select Dividend ETF IDV 11/06/2007 Equity
+iShares Core U.S. REIT ETF USRT 01/05/2007 Real Estate
+iShares Residential and Multisector Real Estate ETF REZ 01/05/2007 Real Estate
+iShares Mortgage Real Estate ETF REM 01/05/2007 Real Estate
+iShares iBoxx High Yield Corporate Bond ETF HYG 04/04/2007 Fixed Income
+iShares Preferred and Income Securities ETF PFF 26/03/2007 Equity
+iShares MBS ETF MBB 13/03/2007 Fixed Income
+iShares 1-5 Year Investment Grade Corporate Bond ETF IGSB 05/01/2007 Fixed Income
+iShares Short Treasury Bond ETF SHV 05/01/2007 Fixed Income
+iShares 3-7 Year Treasury Bond ETF IEI 05/01/2007 Fixed Income
+iShares 5-10 Year Investment Grade Corporate Bond ETF IGIB 05/01/2007 Fixed Income
+iShares Broad USD Investment Grade Corporate Bond ETF USIG 05/01/2007 Fixed Income
+iShares 10-20 Year Treasury Bond ETF TLH 05/01/2007 Fixed Income
+iShares Intermediate Government/Credit Bond ETF GVI 05/01/2007 Fixed Income
+iShares Government/Credit Bond ETF GBF 05/01/2007 Fixed Income
+iShares MSCI KLD 400 Social ETF DSI 14/11/2006 Equity
+iShares Global Industrials ETF EXI 12/09/2006 Equity
+iShares Global Consumer Staples ETF KXI 12/09/2006 Equity
+iShares Global Consumer Discretionary ETF RXI 12/09/2006 Equity
+iShares Global Materials ETF MXI 12/09/2006 Equity
+iShares Global Utilities ETF JXI 12/09/2006 Equity
+iShares S&P GSCI Commodity-Indexed Trust GSG 10/07/2006 Commodity
+iShares U.S. Aerospace & Defense ETF ITA 01/05/2006 Equity
+iShares U.S. Medical Devices ETF IHI 01/05/2006 Equity
+iShares U.S. Home Construction ETF ITB 01/05/2006 Equity
+iShares U.S. Broker-Dealers & Securities Exchanges ETF IAI 01/05/2006 Equity
+iShares U.S. Regional Banks ETF IAT 01/05/2006 Equity
+iShares U.S. Insurance ETF IAK 01/05/2006 Equity
+iShares U.S. Healthcare Providers ETF IHF 01/05/2006 Equity
+iShares U.S. Pharmaceuticals ETF IHE 01/05/2006 Equity
+iShares U.S. Oil & Gas Exploration & Production ETF IEO 01/05/2006 Equity
+iShares U.S. Oil Equipment & Services ETF IEZ 01/05/2006 Equity
+iShares Silver Trust SLV 21/04/2006 Commodity
+iShares Micro-Cap ETF IWC 12/08/2005 Equity
+iShares MSCI EAFE Value ETF EFV 01/08/2005 Equity
+iShares MSCI EAFE Growth ETF EFG 01/08/2005 Equity
+iShares MSCI USA ESG Select ETF SUSA 24/01/2005 Equity
+iShares Gold Trust IAU 21/01/2005 Commodity
+iShares China Large-Cap ETF FXI 05/10/2004 Equity
+iShares Morningstar Mid-Cap Growth ETF IMCG 28/06/2004 Equity
+iShares Morningstar Growth ETF ILCG 28/06/2004 Equity
+iShares Morningstar U.S. Equity ETF ILCB 28/06/2004 Equity
+iShares Morningstar Mid-Cap ETF IMCB 28/06/2004 Equity
+iShares Morningstar Value ETF ILCV 28/06/2004 Equity
+iShares Morningstar Mid-Cap Value ETF IMCV 28/06/2004 Equity
+iShares Morningstar Small-Cap Growth ETF ISCG 28/06/2004 Equity
+iShares Morningstar Small-Cap Value ETF ISCV 28/06/2004 Equity
+iShares Morningstar Small-Cap ETF ISCB 28/06/2004 Equity
+iShares Core S&P Total U.S. Stock Market ETF ITOT 20/01/2004 Equity
+iShares TIPS Bond ETF TIP 04/12/2003 Fixed Income
+iShares Select Dividend ETF DVY 03/11/2003 Equity
+iShares U.S. Transportation ETF IYT 06/10/2003 Equity
+iShares Core U.S. Aggregate Bond ETF AGG 22/09/2003 Fixed Income
+iShares MSCI Emerging Markets ETF EEM 07/04/2003 Equity
+iShares MSCI South Africa ETF EZA 03/02/2003 Equity
+iShares 20+ Year Treasury Bond ETF TLT 22/07/2002 Fixed Income
+iShares 7-10 Year Treasury Bond ETF IEF 22/07/2002 Fixed Income
+iShares iBoxx Investment Grade Corporate Bond ETF LQD 22/07/2002 Fixed Income
+iShares 1-3 Year Treasury Bond ETF SHY 22/07/2002 Fixed Income
+iShares Global Healthcare ETF IXJ 13/11/2001 Equity
+iShares Global Tech ETF IXN 12/11/2001 Equity
+iShares Global Energy ETF IXC 12/11/2001 Equity
+iShares Global Financials ETF IXG 12/11/2001 Equity
+iShares Global Comm Services ETF IXP 12/11/2001 Equity
+iShares MSCI Pacific ex Japan ETF EPP 25/10/2001 Equity
+iShares Latin America 40 ETF ILF 25/10/2001 Equity
+iShares JPX-Nikkei 400 ETF JPXN 23/10/2001 Equity
+iShares North American Natural Resources ETF IGE 22/10/2001 Equity
+iShares MSCI EAFE ETF EFA 14/08/2001 Equity
+iShares Russell Mid-Cap ETF IWR 17/07/2001 Equity
+iShares Russell Mid-Cap Growth ETF IWP 17/07/2001 Equity
+iShares Russell Mid-Cap Value ETF IWS 17/07/2001 Equity
+iShares Semiconductor ETF SOXX 10/07/2001 Equity
+iShares Expanded Tech-Software Sector ETF IGV 10/07/2001 Equity
+iShares U.S. Digital Infrastructure and Real Estate ETF IDGT 10/07/2001 Equity
+iShares Expanded Tech Sector ETF IGM 13/03/2001 Equity
+iShares Biotechnology ETF IBB 05/02/2001 Equity
+iShares Cohen & Steers REIT ETF ICF 29/01/2001 Real Estate
+iShares Global 100 ETF IOO 05/12/2000 Equity
+iShares S&P 100 ETF OEF 23/10/2000 Equity
+iShares MSCI Eurozone ETF EZU 25/07/2000 Equity
+iShares Europe ETF IEV 25/07/2000 Equity
+iShares Core S&P U.S. Growth ETF IUSG 24/07/2000 Equity
+iShares Core S&P U.S. Value ETF IUSV 24/07/2000 Equity
+iShares Russell 2000 Growth ETF IWO 24/07/2000 Equity
+iShares Russell 2000 Value ETF IWN 24/07/2000 Equity
+iShares S&P Mid-Cap 400 Growth ETF IJK 24/07/2000 Equity
+iShares S&P Mid-Cap 400 Value ETF IJJ 24/07/2000 Equity
+iShares S&P Small-Cap 600 Value ETF IJS 24/07/2000 Equity
+iShares S&P Small-Cap 600 Growth ETF IJT 24/07/2000 Equity
+iShares MSCI Brazil ETF EWZ 10/07/2000 Equity
+iShares MSCI Taiwan ETF EWT 20/06/2000 Equity
+iShares U.S. Real Estate ETF IYR 12/06/2000 Real Estate
+iShares U.S. Healthcare ETF IYH 12/06/2000 Equity
+iShares Dow Jones U.S. ETF IYY 12/06/2000 Equity
+iShares U.S. Consumer Discretionary ETF IYC 12/06/2000 Equity
+iShares U.S. Financial Services ETF IYG 12/06/2000 Equity
+iShares U.S. Industrials ETF IYJ 12/06/2000 Equity
+iShares U.S. Energy ETF IYE 12/06/2000 Equity
+iShares U.S. Consumer Staples ETF IYK 12/06/2000 Equity
+iShares U.S. Utilities ETF IDU 12/06/2000 Equity
+iShares U.S. Basic Materials ETF IYM 12/06/2000 Equity
+iShares Russell 1000 Growth ETF IWF 22/05/2000 Equity
+iShares Core S&P Mid-Cap ETF IJH 22/05/2000 Equity
+iShares Core S&P Small-Cap ETF IJR 22/05/2000 Equity
+iShares Russell 2000 ETF IWM 22/05/2000 Equity
+iShares Russell 1000 Value ETF IWD 22/05/2000 Equity
+iShares S&P 500 Growth ETF IVW 22/05/2000 Equity
+iShares S&P 500 Value ETF IVE 22/05/2000 Equity
+iShares Russell 3000 ETF IWV 22/05/2000 Equity
+iShares U.S. Financials ETF IYF 22/05/2000 Equity
+iShares U.S. Telecommunications ETF IYZ 22/05/2000 Equity
+iShares Core S&P 500 ETF IVV 15/05/2000 Equity
+iShares Russell 1000 ETF IWB 15/05/2000 Equity
+iShares U.S. Technology ETF IYW 15/05/2000 Equity
+iShares MSCI South Korea ETF EWY 09/05/2000 Equity
+iShares MSCI Japan ETF EWJ 12/03/1996 Equity
+iShares MSCI United Kingdom ETF EWU 12/03/1996 Equity
+iShares MSCI Canada ETF EWC 12/03/1996 Equity
+iShares MSCI Australia ETF EWA 12/03/1996 Equity
+iShares MSCI Mexico ETF EWW 12/03/1996 Equity
+iShares MSCI Switzerland ETF EWL 12/03/1996 Equity
+iShares MSCI Germany ETF EWG 12/03/1996 Equity
+iShares MSCI Spain ETF EWP 12/03/1996 Equity
+iShares MSCI Singapore ETF EWS 12/03/1996 Equity
+iShares MSCI Hong Kong ETF EWH 12/03/1996 Equity
+iShares MSCI France ETF EWQ 12/03/1996 Equity
+iShares MSCI Italy ETF EWI 12/03/1996 Equity
+iShares MSCI Sweden ETF EWD 12/03/1996 Equity
+iShares MSCI Malaysia ETF EWM 12/03/1996 Equity
+iShares MSCI Netherlands ETF EWN 12/03/1996 Equity
+iShares MSCI Austria ETF EWO 12/03/1996 Equity
+iShares MSCI Belgium ETF EWK 12/03/1996 Equity
+A.4 Extended Tables
 Table A3: Average Euclidean ($\mathcal{L}^{E}$) and Frobenius ($\mathcal{L}^{F}$) distances (×$\times 10^{5}$) using excess returns with different horizons ($F$) for different testing periods: Overall (1 January 2021 to 31 December 2023), Bull-1 (1 January 2021 to 2 January 2022), Bear (3 January 2022 to 12 June 2022), and Bull-2 (13 June 2022 to 31 December 2023).
@@
 $EWMA$ 58.851 90.253 42.125 64.612 94.704 145.253 59.272 90.889
+$PCA$ 71.266 109.164 54.984 83.752 110.811 170.301 70.323 107.864
+$RMT$ 71.064 108.951 54.819 83.588 110.623 170.011 70.094 107.641
 $LW$ 63.548 97.681 47.376 72.757 104.804 160.926 62.036 95.449
@@
 $DCC^{NL}$ 65.361 99.154 51.461 77.640 100.48 152.649 64.160 97.548
-$PCA$ 71.266 109.164 54.984 83.752 110.811 170.301 70.323 107.864
-$RMT$ 71.064 108.951 54.819 83.588 110.623 170.011 70.094 107.641
 $CAB$ 56.954 87.317 44.134 67.401 94.689 145.287 54.291 83.369
@@
 $EWMA$ 44.734 68.528 26.175 40.546 74.009 112.405 48.262 73.923
+$PCA$ 46.390 71.007 24.934 39.238 72.294 109.567 52.781 80.410
+$RMT$ 46.177 70.771 24.796 39.028 72.411 109.269 52.424 80.176
 $LW$ 43.441 66.728 23.834 37.275 76.162 114.702 46.645 71.886
@@
 $DCC^{NL}$ 43.764 67.604 28.135 44.201 75.173 114.919 44.767 69.024
-$PCA$ 46.390 71.007 24.934 39.238 72.294 109.567 52.781 80.410
-$RMT$ 46.177 70.771 24.796 39.028 72.411 109.269 52.424 80.176
 $CAB$ 22.851 35.411 13.091 20.268 42.372 65.754 23.505 36.408
@@
 $EWMA$ 44.704 68.490 26.626 41.728 74.557 112.995 47.751 72.910
+$PCA$ 44.550 68.801 26.466 43.142 72.705 111.010 48.096 73.173
+$RMT$ 44.334 68.575 26.291 42.936 72.803 110.768 47.761 72.940
 $LW$ 42.035 65.084 24.836 40.385 75.018 113.948 43.600 66.895
@@
 $DCC^{NL}$ 42.268 65.595 27.106 43.443 71.582 109.283 43.579 67.259
-$PCA$ 44.550 68.801 26.466 43.142 72.705 111.010 48.096 73.173
-$RMT$ 44.334 68.575 26.291 42.936 72.803 110.768 47.761 72.940
 $CAB$ 22.305 34.990 12.842 20.226 39.166 62.879 23.540 36.458
@@
 $EWMA$ 44.720 68.781 28.925 45.742 73.402 111.890 46.624 71.190
-$LW$ 42.446 66.204 28.694 47.283 79.704 121.665 40.527 62.340
+$PCA$ 44.464 69.171 30.520 50.091 76.867 118.267 44.085 67.265
+$RMT$ 44.239 68.951 30.343 49.925 77.105 118.060 43.693 67.005
+$F=90$ $LW$ 42.446 66.204 28.694 47.283 79.704 121.665 40.527 62.340
 $LW^{F}$ 60.006 91.073 61.704 92.289 64.055 100.713 57.724 87.477
 $CCC$ 43.390 67.342 32.735 51.757 68.495 105.076 42.999 66.475
-$F=90$ $DCC$ 40.234 63.342 29.961 48.303 65.688 101.473 39.492 62.005
+$DCC$ 40.234 63.342 29.961 48.303 65.688 101.473 39.492 62.005
 $DCC^{NL}$ 42.615 66.462 29.762 48.148 71.380 108.855 42.584 66.011
-$PCA$ 44.464 69.171 30.520 50.091 76.867 118.267 44.085 67.265
-$RMT$ 44.239 68.951 30.343 49.925 77.105 118.060 43.693 67.005
 $CAB$ 13.977 21.752 10.080 16.184 20.742 32.133 14.538 22.346
@@
 $EWMA$ 45.330 69.861 33.556 53.195 71.953 109.497 45.222 69.142
+$PCA$ 49.115 76.015 36.715 59.374 84.740 129.651 46.794 71.202
+$RMT$ 48.867 75.827 36.519 59.220 85.041 129.486 46.352 70.985
 $LW$ 47.098 73.175 34.956 56.786 87.296 132.635 43.277 66.502
@@
 $DCC^{NL}$ 44.082 68.671 34.120 54.937 74.174 112.239 41.789 64.902
-$PCA$ 49.115 76.015 36.715 59.374 84.740 129.651 46.794 71.202
-$RMT$ 48.867 75.827 36.519 59.220 85.041 129.486 46.352 70.985
 $CAB$ 12.635 19.816 10.718 17.319 18.659 29.190 12.125 18.707
@@
 $EWMA$ 50.035 77.019 43.437 68.307 76.903 116.122 46.487 71.288
+$PCA$ 58.026 89.446 50.494 79.976 92.884 141.079 52.765 80.559
+$RMT$ 57.744 89.294 50.351 79.843 93.189 140.964 52.223 80.384
 $LW$ 56.137 86.937 49.005 77.757 94.604 142.932 49.567 76.591
@@
 $DCC^{NL}$ 46.532 72.386 41.220 65.248 75.950 114.312 41.414 64.812
-$PCA$ 58.026 89.446 50.494 79.976 92.884 141.079 52.765 80.559
-$RMT$ 57.744 89.294 50.351 79.843 93.189 140.964 52.223 80.384
 $CAB$ 12.356 19.146 11.116 17.602 18.159 27.468 11.472 17.726
@@
 $NA^{F}$ 34.002 52.367 33.076 51.123 64.126 97.309 25.830 40.087
+$PCA$ 77.628 118.284 108.476 165.211 79.544 120.791 57.035 87.075
+$RMT$ 77.315 118.139 108.26 165.074 79.841 120.668 56.481 86.919
 $EWMA$ 53.737 82.720 53.030 82.554 72.648 109.762 48.689 74.953
@@
 $DCC^{NL}$ 40.223 66.604 49.514 77.717 81.610 120.697 22.135 43.633
-$PCA$ 77.628 118.284 108.476 165.211 79.544 120.791 57.035 87.075
-$RMT$ 77.315 118.139 108.26 165.074 79.841 120.668 56.481 86.919
 $CAB$ 16.116 24.758 22.511 34.543 16.201 24.572 11.938 18.457
@@
 $EWMA$ 58.908 90.338 42.125 64.612 94.704 145.253 59.384 91.053
+$PCA$ 71.265 109.164 54.983 83.750 110.813 170.304 70.322 107.863
+$RMT$ 71.064 108.951 54.818 83.587 110.625 170.014 70.093 107.640
 $LW$ 63.547 97.680 47.376 72.755 104.804 160.927 62.035 95.448
@@
 $DCC^{NL}$ 65.338 99.111 51.367 77.516 100.447 152.608 64.186 97.557
-$PCA$ 71.265 109.164 54.983 83.750 110.813 170.304 70.322 107.863
-$RMT$ 71.064 108.951 54.818 83.587 110.625 170.014 70.093 107.640
 $CAB$ 58.272 89.475 43.967 67.109 94.554 145.032 56.996 87.822
@@
 $EWMA$ 44.796 68.618 26.174 40.545 74.005 112.399 48.383 74.100
+$PCA$ 46.394 71.012 24.934 39.237 72.291 109.562 52.789 80.423
+$RMT$ 46.181 70.777 24.796 39.027 72.408 109.265 52.432 80.188
 $LW$ 43.444 66.733 23.834 37.275 76.161 114.700 46.652 71.896
@@
 $DCC^{NL}$ 43.739 67.556 28.016 44.008 75.278 115.106 44.766 69.003
-$PCA$ 46.394 71.012 24.934 39.237 72.291 109.562 52.789 80.423
-$RMT$ 46.181 70.777 24.796 39.027 72.408 109.265 52.432 80.188
 $CAB$ 26.619 41.417 13.427 20.765 45.197 69.510 29.777 46.649
@@
 $EWMA$ 44.759 68.570 26.625 41.727 74.548 112.981 47.861 73.069
+$PCA$ 44.557 68.811 26.466 43.142 72.705 111.009 48.109 73.194
+$RMT$ 44.341 68.586 26.291 42.935 72.803 110.767 47.775 72.961
 $LW$ 42.043 65.095 24.836 40.385 75.020 113.951 43.615 66.916
@@
 $DCC^{NL}$ 42.240 65.544 27.015 43.295 71.326 108.855 43.658 67.380
-$PCA$ 44.557 68.811 26.466 43.142 72.705 111.009 48.109 73.194
-$RMT$ 44.341 68.586 26.291 42.935 72.803 110.767 47.775 72.961
 $CAB$ 24.079 37.328 13.698 21.649 43.977 68.513 25.026 38.428
@@
 $EWMA$ 44.761 68.841 28.925 45.742 73.405 111.893 46.704 71.306
-$LW$ 42.457 66.219 28.694 47.283 79.715 121.682 40.544 62.365
+$PCA$ 44.474 69.186 30.520 50.092 76.878 118.284 44.100 67.288
+$RMT$ 44.249 68.966 30.343 49.925 77.116 118.077 43.709 67.029
+$F=90$ $LW$ 42.457 66.219 28.694 47.283 79.715 121.682 40.544 62.365
 $LW^{F}$ 60.007 91.075 61.690 92.268 64.073 100.740 57.731 87.486
@@
 $DCC^{NL}$ 42.532 66.357 29.680 48.045 70.880 108.147 42.622 66.080
-$PCA$ 44.474 69.186 30.520 50.092 76.878 118.284 44.100 67.288
-$RMT$ 44.249 68.966 30.343 49.925 77.116 118.077 43.709 67.029
 $CAB$ 14.661 23.076 10.115 16.203 22.040 34.086 15.465 24.333
@@
 $EWMA$ 45.361 69.908 33.555 53.194 71.959 109.504 45.282 69.232
+$PCA$ 49.129 76.034 36.715 59.373 84.764 129.686 46.813 71.230
+$RMT$ 48.881 75.846 36.519 59.219 85.065 129.521 46.371 71.013
 $LW$ 47.111 73.194 34.955 56.785 87.320 132.670 43.296 66.530
@@
 $CCC$ 44.126 68.610 36.335 57.587 71.120 108.182 41.325 64.245
-$F=120$ $DCC$ 43.988 68.569 34.032 54.862 73.636 111.512 41.819 64.965
+$DCC$ 43.988 68.569 34.032 54.862 73.636 111.512 41.819 64.965
 $DCC^{NL}$ 44.025 68.619 34.051 54.887 73.690 111.582 41.864 65.025
-$PCA$ 49.129 76.034 36.715 59.373 84.764 129.686 46.813 71.230
-$RMT$ 48.881 75.846 36.519 59.219 85.065 129.521 46.371 71.013
 $CAB$ 12.118 18.831 9.496 15.248 17.864 27.435 12.147 18.653
@@
 $EWMA$ 50.062 77.060 43.438 68.309 76.943 116.140 46.536 71.363
+$PCA$ 58.045 89.474 50.496 79.979 92.932 141.148 52.787 80.592
+$RMT$ 57.763 89.322 50.352 79.846 93.237 141.034 52.245 80.416
 $LW$ 56.156 86.964 49.007 77.759 94.651 143.001 49.589 76.623
@@
 $DCC^{NL}$ 46.406 72.191 41.077 65.104 75.366 113.267 41.432 64.830
-$PCA$ 58.045 89.474 50.496 79.979 92.932 141.148 52.787 80.592
-$RMT$ 57.763 89.322 50.352 79.846 93.237 141.034 52.245 80.416
 $CAB$ 12.828 19.854 11.180 17.728 20.372 30.851 11.702 18.031
@@
 $EWMA$ 53.792 82.799 53.034 82.561 72.673 109.796 48.785 75.090
+$PCA$ 77.678 118.355 108.475 165.210 79.607 120.884 57.115 87.187
+$RMT$ 77.365 118.210 108.260 165.073 79.904 120.760 56.560 87.032
 $LW$ 74.952 114.767 103.983 158.985 80.888 122.122 54.369 83.906
@@
 $DCC^{NL}$ 40.261 66.455 49.351 77.363 80.796 119.188 22.552 44.012
-$PCA$ 77.678 118.355 108.475 165.210 79.607 120.884 57.115 87.187
-$RMT$ 77.365 118.210 108.260 165.073 79.904 120.760 56.560 87.032
 $CAB$ 18.246 27.828 29.049 44.148 14.976 22.681 12.182 18.727
-Table A4: Performance of the GMV portfolios constructed using the covariance forecasts from the 11 models under consideration plus the equal-weighted portfolio for the different market regimes.
-The portfolios are compared based on their annualised out-of-sample standard deviation ($\sigma_{m}$) and average out-of-sample turnover ($\text{TO}_{m}$), respectively.
-Daily Weekly Monthly
-$\sigma_{m}^{2}$ $TO_{m}$ $\sigma_{m}^{2}$ $TO_{m}$ $\sigma_{m}^{2}$ $TO_{m}$
-Bull-1 $1/N$ 0.009883 0.005169 0.009873 0.002335 0.009907 0.001103
-$NA$ 0.000543 0.150112 0.000794 0.060372 0.000834 0.026885
-$NA^{F}$ 0.000731 0.001378 0.000732 0.000607 0.000736 0.000327
-$EWMA$ 0.000735 0.088541 0.000748 0.043884 0.000754 0.015613
-$LW$ 0.001054 0.091979 0.001572 0.044883 0.001692 0.022283
-$LW^{F}$ 0.000780 0.001228 0.000781 0.000610 0.000783 0.000349
-$CCC$ 0.000620 0.119486 0.000681 0.053221 0.000728 0.013605
-$DCC$ 0.000529 0.207262 0.000698 0.093558 0.000781 0.024324
-$DCC^{NL}$ 0.000522 0.217432 0.000784 0.097809 0.000875 0.034673
-$PCA$ 0.000850 0.236835 0.000827 0.090803 0.000795 0.020906
-$RMT$ 0.000856 0.150601 0.000860 0.067402 0.000826 0.025859
-$CAB$ 0.000634 0.086489 0.000656 0.055931 0.000707 0.023736
-Bear $1/N$ 0.029939 0.007221 0.029940 0.003443 0.029793 0.001599
-$NA$ 0.002207 0.092111 0.002682 0.032018 0.002959 0.018142
-$NA^{F}$ 0.002891 0.001650 0.002892 0.000935 0.002906 0.000436
-$EWMA$ 0.002908 0.079065 0.002843 0.036314 0.003120 0.019997
-$LW$ 0.003118 0.095723 0.003970 0.041663 0.004754 0.021610
-$LW^{F}$ 0.003196 0.002035 0.003197 0.001043 0.003201 0.000570
-$CCC$ 0.002497 0.129487 0.002646 0.043843 0.002590 0.012451
-$DCC$ 0.002355 0.067768 0.002639 0.015372 0.002636 0.005911
-$DCC^{NL}$ 0.002269 0.083567 0.002652 0.024277 0.002730 0.010308
-$PCA$ 0.002800 0.114256 0.002792 0.043637 0.002864 0.011111
-$RMT$ 0.002983 0.094396 0.002896 0.037508 0.003153 0.023308
-$CAB$ 0.002273 0.030506 0.002464 0.022686 0.002656 0.016836
-Bull-2 $1/N$ 0.019967 0.005666 0.019956 0.002382 0.019933 0.001334
-$NA$ 0.002823 0.118997 0.003526 0.053457 0.003667 0.029686
-$NA^{F}$ 0.003398 0.001426 0.003400 0.000577 0.003406 0.000261
-$EWMA$ 0.003432 0.073812 0.003433 0.034460 0.003630 0.016887
-$LW$ 0.003512 0.102531 0.004274 0.048313 0.004245 0.020765
-$LW^{F}$ 0.003874 0.002113 0.003879 0.001123 0.003889 0.000788
-$CCC$ 0.003233 0.129874 0.003635 0.047812 0.003671 0.015276
-$DCC$ 0.002947 0.071444 0.003294 0.026871 0.003364 0.009476
-$DCC^{NL}$ 0.002924 0.085242 0.003318 0.034070 0.003365 0.012661
-$PCA$ 0.003793 0.194898 0.003678 0.063899 0.003597 0.024379
-$RMT$ 0.003668 0.118539 0.003720 0.055498 0.003762 0.029837
-$CAB$ 0.003323 0.087821 0.003547 0.052596 0.003885 0.029487